### v1

# Explainable Metrics for the Assessment of Neurodegenerative Diseases through Handwriting Analysis

Abstract
Motor changes are early signs of neurodegenerative diseases (NDs) such as Parkinson’s disease (PD) and Alzheimer’s disease (AD), but are often difficult to detect, especially in the early stages. In this work, we examine the behavior of a wide array of explainable metrics extracted from the handwriting signals of 113 subjects performing multiple tasks on a digital tablet. The aim is to measure their effectiveness in characterizing and assessing multiple NDs, including AD and PD. To this end, task-agnostic and task-specific metrics are extracted from 14 distinct tasks. Subsequently, through statistical analysis and a series of classification experiments, we investigate which metrics provide greater discriminative power between NDs and healthy controls and among different NDs. Preliminary results indicate that the various tasks at hand can all be effectively leveraged to distinguish between the considered set of NDs, specifically by measuring the stability, the speed of writing, the time spent not writing, and the pressure variations between groups from our handcrafted explainable metrics, which shows $p$-values lower than 0.0001 for multiple tasks. Using various classification algorithms on the computed metrics, we obtain up to 87% accuracy to discriminate AD and healthy controls (CTL), and up to 69% for PD vs CTL.
1 Introduction
Neurodegenerative diseases (NDs), such as Parkinson’s disease (PD) and Alzheimer’s disease (AD), have detrimental effects on specific regions of the brain, resulting in a gradual deterioration of cognitive, functional, and behavioral capabilities. This deterioration can be monitored, which is why handwriting, with its complex nature that involves cognitive, kinetic, and perceptual motor components, plays a significant role in the evaluation of NDs [1, 2]. Handwriting involves crucial brain regions related to learning and execution functions, including the cerebral cortex, basal ganglia, and cerebellum [3], which is why alterations in its patterns can serve as source of prominent biomarkers, indicative of both the presence and severity of the disease [4].
Paper and digital handwriting tests [4, 5, 1] have been used to assess the presence of NDs through various signs. Previous works have been mainly focusing on handwriting for PD assessment [6, 7, 8, 9], with a handful starting to tackle AD assessment as well [10]. Subjects with PD often suffer from micrographia [11, 12], tremor [12], bradykinesia [12], and hypometria [13], a collection of signs collectively referred to as PD dysgraphia[1].
The handwriting tests used for the possible assessment of NDs can be broken down into several categories that are used to collect various metrics. The first category is simple drawing tasks, including spirals and straight lines. Simple drawings can be used to assess trajectory, tremor, dimensionality, velocity, and acceleration [14]. The second category is simple writing tasks, including simple words and short sentences. When performed on a digital tablet, these tasks allow for the analysis of in-air movements and motor planning between words [14]. Collecting in-air movements is important because the time in-air results from a person planning out their next movement, which reflects their cognitive ability [5]. Additionally, in-air and on-surface movements were found to be statistically independent and can provide non-redundant information about the writer [15]. The in-air time is also indicative of hesitation and motor planning activities. Fluid handwriting, or handwriting with minimal in-air time, demonstrates the presence of an anticipated motor plan. At the same time, longer hesitations between writing might show that the writer needs to rethink their motor plan. The third category is complex tasks, which include complex drawing or paragraph copying. Complex drawing tasks and freewriting, as in the clock drawing task [16], allow for the detection of visual-spatial deficits by analyzing a person’s spatial organization skills. Copying tasks provide comparisons of writing variations with respect to different stimuli [17]. The use of a variety of these tasks provides a full profile of a person’s handwriting capabilities.
In this study, we collected a new data set from 103 participants with various NDs, such as AD, Mild Cognitive Impairment (MCI) due to AD, PD, and various Parkinsonisms that are difficult to differentiate from PD and that we will name Parkinson’s Disease Mimics (PDM). This dataset comprises recordings of handwriting on a digital tablet across 14 specific tasks designed to assess neurodegenerative conditions. We extract explainable and quantitative endpoints, also called metrics, from 14 handwriting tasks and pinpoint key metrics capable of distinguishing between healthy controls and various NDs groups, including an active control group to ocmpare to PD participants: the PDM. We first explore a more general approach to obtain task-agnostic metrics, independently of their original purpose or content.
Secondly, a task-dependant approach is adopted, to fully explore which traits of the participants can be used for discrimination, following the expected behavior that each group of participant is supposed to exhibit for each group of tasks.
Furthermore, we use statistical analysis and machine learning classifiers to compare the prediction power of the extracted metrics, using binary comparisons between groups. The results obtained from that analysis highlight the possible value of these endpoints as digital biomarkers.
The main contributions of this article are:
  • The introduction of a handwriting dataset for the assessment of NDs
  • The inclusion of several ND groups, including an active control group to compare to PD (the PDM group), where all precedent studies were limited to studying one or two NDs.
  • The use of Handwriting for AD and MCI assessment, usually reserved to PD assessment.
  • A large set of task-agnostic and task-specific metrics that show significant differences between groups of participants, following expected behaviors.
  • The introduction to inter-tasks comparisons of metrics, to study the behavior differences for a given participants facing a difference in their goals, such as changing the drawing hand.
The following sections of the manuscripts are organized as follows:
  • Section 2 presents the data collection pipeline and the main pre-processing stages.
  • Section 3 introduces the task-agnostic metrics extracted from all tasks and the task-specific metrics designed for specific groups of tasks. Then, various methods used to gauge the significance of our metrics in the various tasks analyzed are described. Finally, we introduce the methods used for training classifiers for NDs assessment from the presented explainble metrics.
  • Section 4 shows and discusses our statistical and classification results.
  • Section 5 contains the main findings and conclusions of the study.
2 Materials
In this section, we outline the data collection and preprocessing methods, along with the tasks evaluated in this research.
2.1 Data Collection
The authors of this study collected a dataset from 113 participants [18, 19, 20, 21, 22, 23]. These participants were either categorized as healthy (referred to as CTL) or diagnosed and treated for a neurodegenerative disease by a specialist neurologist or geriatrician from Johns Hopkins Hospital. Participants diagnosed with NDs were categorized into one of several groups based on their specific diagnosis:
  • AD participants clinically diagnosed and treated by an AD expert geriatrician.
  • MCI participants presenting signs of MCI, evaluated by the same geriatrician, but without any AD diagnosis yet.
  • PD participants with clinically established diagnosis and being treated by a movement disorders neurologist.
  • PDM participants presenting symptoms similar to those of PD, evaluated by the same neurologist, but with a diagnosis different from PD. This group includes patients initially diagnosed with Dementia with Lew Bodies, Fahr’s Disease, Ataxia, Corticobasal Syndrome, Corticobasal Degeneration, Cervical Dystonia, Gerstmann Schenker Schyuler Syndrome, Wilson Disease, MSA-Parkinsonian, and Essential Tremor. In this study, this group is considered an active control group against the PD group.
All participants signed informed consent, and the data collection was approved by the Johns Hopkins Medical Institutional Review Board. Participants with PD continued their usual pharmacological treatment and took dopaminergic medication before the recording session. Handwriting samples were collected using a Wacom One 13 tablet with the associated pen. This setup enabled us to record the pressure applied by the pen on the tablet surface, as well as its position both on the tablet and in the air when in close proximity. Additionally, the recordings encompass the synchronized speech of participants and their ocular eye movements, although these two modalities were not utilized in this study. Participants completed 14 written tasks per session, detailed in Section 2.2, guided by a research assistant. We collected recordings for each participant for at least one session, while some participants were recorded up to three times, when available and interested, with each session spaced six months appart. The distribution of participants, with their age, gender, and total number of tasks recorded, is presented in Table 1 Participants classified as AD and MCI were combined into a single group, named AD in this study. Only a subset of the recorded participants were kept in that study, to balance as much as possible the age and gender of the participants. However, gender balancing for the AD group was not possible, given the low number of female participants, so we choose to keep the whole group.
Table 1: Distribution of the participants recorded in the dataset.
Category Number of % Females Age Files MoCA UPDRS3
Participants recorded
CTL 42 57.1% 69y (±$\pm 11$) 757 25.4 17.0
PD 35 40.0% 67y (±$\pm 9$) 608 25.9 24.3
PDM 15 46.7% 53y (±$\pm 14$) 283 25.8 21.5
AD 21 19.0% 70y (±$\pm 7$) 745 19.7
Total 113 43.4% 66y (±$\pm 12$) 1840
2.2 Handwriting Tasks
In each session, we collected data from 14 handwriting tasks per participant. Due to various reasons, technical or human, the recordings of some tasks can fail, leading to an different number of participants for each task. Participants having one or more missing tasks are still included in the study. Within this section, we provide a comprehensive description of tasks, their origin (what each of those tasks were used for, originally), and elucidate the experimental hypotheses we formulated for each. Tasks such as Point and Spiral tasks involved participants performing stability exercises three times: alternating hands, then with their dominant hand while concurrently speaking. Writing and Drawing tasks prompted participants to replicate an item from the screen, from memory, and generate new content. Participants were asked to not touch the tablet with their arm or hand while writing. Figure 1 illustrates a sample from each task type, encompassing participants from the CTL, AD, PD, and PDM groups.
Figure 1: Example of each type of task from all the four experimental groups. When on tablet, the color of the points transits from red to purple. When in air, the points are depicted in gray.
2.2.1 Point Tasks
[…]
2.2.2 Spiral Tasks
The spiral tasks are also designed to assess the motor capacities of the participants. Multiple studies require participants with PD to perform regular circular movement on paper or on tablet to assess their regularity or the presence of an eventual tremor [24, 25, 26]. The spiral test has also proven useful for AD assessment using a tablet [27]. The participants were asked to draw a spiral starting from the center of the tablet. More specifically, in the Spiral Dominant and Spiral Non-Dominant tasks, participants were asked to use their dominant and Non-Dominant hand respectively, and in the Spiral PaTaKa, they were asked to use their dominant hand, in addition, to perform a diadochokinetic task, i.e., repeat as fast as possible the syllables pa, ta, and ka in this exact order. However, our current focus lies solely on the handwriting data. We hypothesize that the simultaneous performance of tasks will make it more difficult for participants to use strategies to compensate for the motor impairments caused by NDs.
2.2.3 Writing Tasks
The writing tasks were designed to assess the cognitive and motor capacities of the participants. Handwriting analysis has been used mostly for the assessment of PD, as some motor symptoms of the disease can be seen through some characteristic modifications of the handwriting, such as micrographia [11, 28]
Participants first engaged in the Copy Text task, wherein they were tasked with replicating a paragraph of text displayed on the screen. We then focus on the impact of multitasking by introducing the Copy Read Text task, where participants copied another text while simultaneously reading it out loud. Subsequently, participants’ spontaneous writing abilities were captured through the Freewrite task, prompting them to write freely one or two sentences that must not contain any personal information. Lastly, a different cognitive function was explored via the Numbers task, wherein participants were tasked with solving eight simple arithmetic operations involving the sum of two natural numbers not exceeding a total of 20, while saying the results of the sums out loud. It is worth noting that the Numbers task may be analyzed differently from the writing tasks due to its repetitive nature, necessitating a distinct analytical approach.
2.2.4 Drawing Tasks
The Drawing tasks are also designed to assess the cognitive and motor capacities of the participants.
First, in the Copy Image task, we asked participants to copy an image displayed on the screen. The image is shown in the figure 2
Figure 2: Target image presented to the participants for the copy image and copy image memory tasks.
We further examined the impact of memory on this behavior by introducing the Copy Image Memory task. After copying the image in the Copy Image task, participants were then asked to reproduce it from memory approximately five minutes later.
Subsequently, participants were asked to perform two tasks: First, the Draw Clock task, which looked at their proficiency in drawing all the required elements of a clock at eleven past ten, used for AD assessment in the MoCA test [29] and for PD assessment [16], and for discrimination between neuro-degenerative diseases [30].
Then, the Draw Cube task, used in the MoCA test [29] and which has been proved to be also useful for PD assessment [31], examines their capacity to represent a three-dimensional object.
2.3 Data Curation
Every task starts and concludes with a participant crossing an area marked as start or end on the screen, respectively, using a pen. However, the portion of the recording relevant to the analysis is shorter than the entirety captured. Hence, we manually eliminated the empty segments before and after the tasks commenced. Files lacking task-related responses are removed.
3 Methods
This section describes the metrics we computed for each task and the approach used to assess their significance between the various experimental groups considered. We meticulously crafted a set of interpretable metrics derived from the raw handwriting signals collected during the various tasks of interest. Our analysis of the various tasks is twofold: firstly, we scrutinize the dynamic properties of handwriting, which are common to all tasks. Secondly, we evaluate their performance in specific tasks, such as stability for point tasks and proximity between drawing and target for copying tasks. Dynamic analysis is conducted across all relevant tasks to extract overarching metrics concerning handwriting, which will be used to measure certain behavioral aspects of each disease. Conversely, task-specific metrics provide deeper insights into particular aspects of the disease by investigation of a limited set of tasks at a time.
3.1 Metrics computation pipeline
To compute the metrics analyzed in this study, we follow the pipeline described in Figure 3. Each data file can be described as a time series containing $N\in\mathbb{R}$ points:
  • The positions $X=(x_{i})_{i\in[[1,N]]}$ and $Y=(y_{i})_{i\in[[1,N]]}$
  • The time stamps $T=(t_{i})_{i\in[[1,N]]}$
  • The pressure $P=(p_{i})_{i\in[[1,N]]}$
We define a clean file as the list $\mathcal{F}=[X_{u},Y_{u},P_{u},T_{u}]$.
We computed various metrics from the clean files, targeting various cognitive and behavioral elements. The metrics can be categorized into two distinct sets:
  1. 1.
    Task-agnostic (see Section 3.2): these metrics measure primarily dynamic properties of the signal such as speed, noise, and basic ratios.
  2. 2.
    Task-specific (see Section 3.3): these metrics are designed for specific tasks or groups of tasks.
3.2 Task-Agnostic Metrics
Task-agnostic metrics quantify the dynamic aspects of the drawings. They are computed using various signal-processing techniques. The extraction process, described in Figure 3, consists of three distinct steps:
  1. 1.
    computation of various time series from the raw data, such as the angle between 2 consecutive points, the distance between two points, and multiple derivations (see Section 3.2.1).
  2. 2.
    segmentation of each series using only the segments gathered In-Air or On-Tablet, and additionally decomposing the On-Tablet sequence in a suite of strokes. The segmentation is described in Section 3.2.2
  3. 3.
    computation of statistics (e.g., mean, the standard deviation) on the time-series obtained from the derivation of time independent metrics, described in Section 3.2.3.
Figure 3: Pipeline used to extract task-agnostic metrics across all tasks.
3.2.1 Time series
For every point $i\in[[2,N]]$ of a given clean data file $\mathcal{F}$, we compute:
  • The angle between 2 consecutive points $\alpha_{i}=arctan(y_{i}-y_{i-1},x_{i}-x_{i-1})$.
  • The $L2$ norm between 2 consecutive points $n_{i}=\sqrt{(x_{i}-x_{i-1})^{2}+(y_{i}-y_{i-1})^{2}}$.
Subsequently, we calculate the first and second derivatives for the following:
  1. 1.
    Angle: angular velocity and angular acceleration.
  2. 2.
    Norm: velocity and acceleration.
  3. 3.
    Positions: horizontal and vertical speed and acceleration.
  4. 4.
    Pressure: pressure variations.
These computations yield a total of 10 columns of data, which are utilized in subsequent analyses.
3.2.2 Segmentation
We implemented segmentation to distinguish between segments occurring on the tablet surface and those in the air. Subsequently, we computed time-independent metrics for the entire file, focusing either solely on points recorded on the tablet or those recorded in the air. Given that the point task primarily involves actions in the air, statistical computations are exclusively performed on the in-air segments. Conversely, for spiral tasks, only segments registered on the tablet are utilized. However, for writing and drawing tasks, both in-air and on-tablet segments are utilized. Notably, metrics based on pressure are not retained for the in-air segments, as they would be derived from a constant value.
3.2.3 Time-Independent Metrics
From the time series computed previously, we extract a range of statistics to characterize the dynamic properties of the signal. The list of statistics includes, for a given signal $S=(s_{i})_{1\leq i\leq N}$:
  • Mean ($\mu):\frac{1}{N}\sum_{i=1}^{N}s_{i}$
  • Standard deviation ($\sigma):\sqrt{\frac{1}{N}\sum_{i=1}^{N}(s_{i}-\mu)^{2}}$
  • Amplitude ($A):max(S)-min(S)$
We chose not to utilize the amplitude measure for angular velocity and accelerations, as it would always yield identical values. Furthermore, we calculated the Shannon Entropy based on the points’ positions, whether on the tablet or in the air. To achieve this, we employ a Gaussian kernel density estimate (KDE) on the points’ $X$ and $Y$ positions. This enabled us to determine the probability $q_{i}$ for each point $(x_{i},y_{i})$. Shannon Entropy is then defined as $H_{s}=-\sum_{i=1}^{N}q_{i}ln(q_{i})$. Following the computation of these metrics, we obtained a total of 54 task-agnostic metrics across the 14 tasks.
3.3 Task-Specific Metrics
In this section, we detail the various task-specific metrics we computed. We divided the presentations of the metrics along the four main tasks’ categories proposed in Section 2.2, with the addition of a section specific to the Numbers task.
3.3.1 Point Tasks Metrics
This section presents the metrics specific to the Point Tasks. Figure 4 depicts these metrics. Point Tasks are mainly used to measure the motor skills of participants and are mostly used for PD assessment. Most dynamic metrics have already been explored as task-agnostic metrics, this is why we focused on the static metrics of the drawing: how symmetric and regular it is, and what the dispersion of the samples is. For this task, we calculated the geometric center of the pen’s position over the digital tablet. We used this geometric center as a pivot to compare the distribution of the points with respect to the computed center, through 3 metrics:
  1. 1.
    Average Radius measures the dispersion of the pen positions related to the geometric center, a first measurement of stability, which measure the amplitude of the variations.
  2. 2.
    Standard deviation of the Radius measures the variations in the distance from the center, a second measure more focus on intermittent tremors, not on regular vibrations.
  3. 3.
    Average Angle measures the average of the angle between a point and the center, which can be used as a proxy for the symmetry of the drawing. Tremor and bradykinesia in PD do not always affect all muscles, if only certain muscles of the arm were affected, we
Figure 4: Schematic of the characteristics measured by different point-tasks specific metrics.
3.3.2 Spiral Tasks Metrics
This section discusses static and dynamic metrics used to analyze hand-drawn spirals.
Those metrics are following three directions, illustrated in Figure 5:
  1. 1.
    Static characteristics of the ellipse bounding the spiral.
  2. 2.
    Static characteristics of the inside loops of the spiral.
  3. 3.
    Dynamic characteristics of the drawing, based on the time taken.
Figure 5: Spiral-specific metrics computation process
Ellipse characteristics
As shown in Figure 5, the participants produce rounded spirals for all the spiral tasks. Our approach involves initially characterizing the properties of the outer boundaries of the spiral as an ellipse, leveraging these properties as metrics. We used Principal Component Analysis (PCA) to find the directions of the semi-major and semi-minor axes of each spiral, then find the points further away from the geometric center of the spiral in both of those directions to obtain a measure of the semi-minor axis $a$ and the semi-major axis $b$. Using those values, we define a set of metrics:
  • Area of the ellipse: $Area=\pi*a*b$.
  • Eccentricity: $Ec=\sqrt{1+\frac{b^{2}}{a^{2}}}$. The eccentricity is a scalar [0,1[$\in$[0,1[∈ [ 0 , 1 [ that measures the compression of an ellipse when defined as a conic section. An eccentricity of 0 would be the one of a circle, while an eccentricity of 1 would indicate a parabola.
  • Ellipticity: $El=\frac{b}{a}$.
  • Degrees of rotation: The angle between the semi-major axis and the abscissa.
Spiral arms
Once the outside shape has been characterized, we looked into the properties of the arms of the spiral. Each spiral is split into segments, containing each one full loop. For each loop, we computed its radius as the average distances between its points and the geometric center of the spiral, which allowed us to compute another set of metrics:
  • Number of loops
  • The Average distance between loops: the difference between 2 loops radii is defined as the distance between those 2 loops. We take the mean of this measure for all pairs of consecutive loops.
  • The standard deviation of the loops radius differences: the standard deviation of the radius difference between consecutive loops. This metric quantifies the regularity of the spiral.
  • The slope of the loops radius differences: the slope of the radius difference between consecutive loops.
Dynamic analysis
Dynamic metrics are then computed based on the total time taken to complete the task:
  • Duration: total time taken to draw the spiral.
  • Time per loop: the average time taken to draw one loop.
  • Frequency: the average number of loops computed per second.
3.3.3 Numbers Task Metrics
This section focuses on the metrics unique to the Numbers task. Participants engaged in solving addition problems by moving down the tablet, allowing for individual analysis of each response. This methodology enabled us to track the evolution of handwriting metrics as participants navigate through the task. For a comprehensive breakdown of these metrics, please refer to Figure 6 below.
Figure 6: Schematic of the metrics used for each answer of the Numbers task. The left graph shows the variations in pressures across the whole task, and the right schematic shows the metrics measured between 2 consecutive answers.
Average Width of Answer
The average width of each answer quantifies the size of the handwriting. It is computed by subtracting the minimum $X$ coordinate from the maximum $X$ coordinate for each answer and then averaging these values. People with PD have been found to have a lower control of their movement amplitudes, resulting in longer strokes and larger handwriting [32].
Change in Width of Answer
The variation of width from one answer to the next. It is computed through a linear regression between the width of each answer and the question number. The metric itself represents the slope of this regression line. The variation in width is aimed to measure the presence of micrographia [11].
Change in Average Answer Pressure
The average answer pressure metric is a measure of the amount of pressure that a person applies while writing the answer. This metric can be leveraged as a time series metric to track its variation as the participant progresses through the task. It is computed through a linear regression between the average pressure of each answer and the question number. The metric itself represents the slope of this regression line.
Time per Answer
The total duration metric measures the time a participant spends answering one of the questions. Given the absence of a predefined time limit for the Numbers task, this metric provides insight into the participant’s pace in completing the task.
3.3.4 Writing Tasks Metrics
In this section, we present the metrics extracted from the writing tasks: freewrite, copy text, and copy text memory[33]. The computation of these metrics is shown in Figure 7.
Figure 7: Schematic of the metrics used in writing tasks, illustrated for the words ’This is an example’.
Pressure Standard Deviation
The pressure standard deviation metric is a measure of the spread of the amount of pressure a person applies while writing. This measures how steady the hand is when writing by capturing the variations in pressure. Only the nonzero pressure values were considered when calculating this metric.
Average Time Between Writing
The average time between writing measures how long it takes for the participant to start writing after finishing a word. This shows how fluent the writing is and how stable the motor planning activity is. The metric was calculated by averaging the total time the pen was not touching the tablet.
Average Horizontal Distance Between Words
This metric was calculated by finding the difference between the $X$ value of when the pen is lifted and when it is put back down. Only the $X$ distance was considered because the $Y$ distance between the words is dependent on both the letters they are writing and how they write the letters.
Total Time
The total time metric gauges the duration required for task completion. It is determined by calculating the time elapsed from when the participant initially places the pen to start writing until they lift the pen after writing the last word. This metric holds greater significance for tasks like Copy Text and Copy Read Text, where participants’ required text to write is the same.
3.4 Statistical Analysis
After extracting a total of 76 metrics (i.e., 54 task-agnostic metrics and 22 task-specific metrics), we measure which can be used to assess and characterize NDs. Our goal is twofold: firstly, to discern which metrics can effectively differentiate between the CTL group and the other experimental groups under consideration, and secondly, to ascertain whether handwriting metrics alone can distinguish PD from other forms of Parkinsonism (PDM). To achieve this, we conduct Kruskal’s H-test across all possible pairs of groups for each computed metric, using the Scipy toolkit [^1] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.kruskal.html. Specifically, we examine the ability of each metric to distinguish the CTL group from each disease group, which requires the following comparisons: AD vs. CTL, PD vs. CTL, and PDM vs. CTL. We also assess the effectiveness of each metric in differentiating the PD from the PDM group. This procedure yields four comparisons, resulting in four p-values per metric.
The underlying experimental hypothesis is that the information in a given task can be leveraged to extract descriptors that can discriminate between two groups. To avoid false discovery, we apply FDR correction (from the Statsmodel toolkit [^2] https://www.statsmodels.org/dev/generated/statsmodels.stats.multitest.fdrcorrection.html on all the p-values grouped by task and test, so we apply $4\times 14=56$ corrections. Significant results from the statistical analysis are presented and commented on in the following section.
To complete our analysis of the separability of those groups of participants, we measured the Area Under the Curve (AUC) for pairs of participant groups, which will be presented alongside FDR-corrected p-values.
3.5 Additional comparisons
The tasks analyzed consist of a carefully selected set of closely related tasks to enable comparative analysis of a specific type of task. Two types of comparisons are proposed:
  • Grouped metrics: Instead of measuring the significance of the metrics task by task, we compare metrics for groups of tasks.
  • Differences of metrics: Instead of considering the metrics, we consider the difference between the same metric evaluated on 2 different tasks.
Figure 8 shows how we perform those operations for the different tasks.
Figure 8: Schematic of the organization of the tasks when comparing 2 tasks or grouping a set of tasks.
3.5.1 Grouped metrics
We supposed that certain of the metrics extracted for a given type of task should have a coherent behavior across all tasks. To verify this, we grouped them: for each group of tasks, the same statistical comparisons as the ones detailed in section 3.4 are performed, but considering all recordings of a group of tasks instead of the recordings of a specific task. We measured the significance of all metrics for each of those task groups as one task:
  • All spiral tasks: Spiral dominant, Spiral non-dominant, and Spiral PaTaKa.
  • All point tasks: Point dominant, Point non-dominant, and Point sustained.
  • All writing tasks: Freewrite, Copy Text, and Copy Read Text.
  • All drawing tasks: Draw Clock, Draw Cube, Copy Image, and Copy Image Memory.
Table 2 summarizes the tasks at hand sorted per group and the specific metrics proposed per group of tasks.
Table 2: summary of the tasks at hand and their specific metrics.
Group of tasks Tasks Description Specific Metrics
Point Right Maintain the pen vertically above a point without the Radius from centroid
Points Point Left moving or touching the tablet, using their right hand, Angle from Centroid
Point Sustained left hand, or dominant hand while sustaining the vowel /a/. Std of the Radius
Area of the spiral
Spiral Right Eccenctricity
Drawing a round spiral from the center Ellipticity
Spirals Spiral Left using their right hand, left hand, Number of loops
or dominant hand while repeating PaTaKa. Distance between loops
Spiral PaTaKa Time per loop
Frequency
Answer 8 consecutive arithmetic Avg width of answer
Numbers - questions and writing the Change in Avg pressure
answer while saying it. Time per answer
Pressure Ratio
Copy Text Copy a small paragraph being shown, Pressure std.
Writing Copy Text Memory then copy the same from memory, Avg time between words
Freewrite finally writing a freeform paragraph. Avg distance between words
Total time
Copy Image Copy an image being shown,
Drawing Copy Image Memory then copy the same from memory, CLIP similarity image-image
Draw Clock drawing a clock CLIP similarity text-image
Copy Cube and drawing a cube.
3.5.2 Difference between metrics
To study the variations between tasks and how changes of behavior might be signs of NDs, we compared the metrics extracted between 2 tasks. For each session, if a given participant had successfully completed both tasks of a pair, we took the difference between 2 metrics, and perform the same statistical analysis as described in section 3.4 for this new ’coupled’ metric. The considered couples are the following:
  • Dominant vs. Non-Dominant: The difference when drawing with their non-dominant hand is computed to measure eventual dissymmetry in their response. This difference is measured for the following tasks:
    • Point Dominant vs Non Dominant
    • Spiral Dominant vs Non Dominant
  • Reading vs Non-Reading: The difference when performing multiple tasks at the same time is measured by comparing the following tasks:
    • Point Dominant vs Point Sustained
    • Spiral Dominant vs Spiral PaTaKa
    • Copy Text vs Copy Read Text
  • Memory vs Non-Memory: The difference when a task from memory or from a given image is measured by comparing the task Copy Image against the task Copy Image Memory.
This gave us 6 new potential comparisons.
3.6 Classification Experiments
In addition to conducting statistical analyses, we conduct binary classification experiments across various experimental groups using metrics extracted from each task. These experiments further evaluate the discriminative power of the metrics and the utility of tasks in extracting valuable information for characterizing different disorders. We are employing a range of classifiers, including Support Vector Machine (SVM), K-Nearest Neighbors (KNN), Random Forest (RF), Extreme Gradient Boosting (XGBoost), Bagging (BG), and Multi-Layer Perceptron (MLP)  [^3] To implement the different classifiers, we are using the Scikit-learn (Sklearn) library in Python. See https://www.tutorialspoint.com/scikit_learn/scikit_learn_introduction.htm.
A Nested Cross-Validation (NCV) scheme is being applied in our experiments. NCV is a common approach that can be adopted for metric selection and parameter tuning to obtain reliable classification accuracy and prevent overfitting [34, 35]. We are splitting the data set into ten outer folds, and each fold is held out for testing while the remaining $k-1$ folds (9 folds) are merged to form the outer training set. Each outer training set is further split into stratified inner folds for inner training and validation. We identify the 30 best metrics (IFMs only) and hyperparameters in the inner folds via grid search. An Extremely Randomized Trees Classifier is used to perform metric selection. [^4] The documentation of the Python library used to implement the Extremely Randomized Trees Classifier can be found at https://scikit-learn.org/stable/modules/generated/sklearn.ensemble.ExtraTreesClassifier.html. It is a type of ensemble learning technique which aggregates the results of multiple de-correlated decision trees to output its classification result. Ten sets of unique best configurations are obtained after getting the best hyperparameter configuration for each outer iteration using their corresponding inner folds.
To improve stability in metric and hyperparameter selection with small datasets, we adapt the NCV process. We determine the final hyperparameter configuration by averaging (or using the mode of) the best configurations from ten iterations. The model is then trained and tested from scratch using this configuration on distinct outer training/test folds, ensuring no overlap between training and testing participants. This approach includes 10-fold cross-validation for hyperparameter tuning with all training data and applies NCV separately for each task. Data is standardized within each outer iteration using the training set’s mean and standard deviation. Figure 9 outlines the classification pipeline and NCV scheme.
Figure 9: a) Machine learning pipeline was adopted for training and testing. b) Nested cross-validation for algorithm selection. In the graph, the nested cross-validation process is exemplified using 3 folds for training, one for testing in the outer loop, 2 for training, and one for validating in the inner loop. However, in our experiments, we used 9 folds for training, one for testing in the outer loop, 8 for training, and one for validating in the inner loop. This strategy was applied separately when training each of the classifiers adopted in the experiments.
4 Results
In this section, we first present the most significant metrics obtained by groups of metrics for each test realized. Then, we comment on the results obtained in the classification experiments for each task.
4.1 Task-agnostic Metrics Results
Out of 54 task-agnostic metrics computed, 53 are significant ($p\leq 0.05$) at least for one task and one comparison between groups. The complete table is available in the Appendix; however, we prioritize limited sets due to the considerable number of significant metrics. We present the metrics most effective in distinguishing AD and PD from control participants and separate PD from PDM participants. Subsequently, we will highlight metrics that exhibited robust behavior across multiple comparisons.
4.1.1 AD vs CTL
Over the 54 task-agnostic metrics, 48 are significant ($p\leq 0.05$) in comparing AD and CTL participants for at least one task, and 17 have p-values smaller than 0.001. We present the 10 metrics having the lowest p-values in Table 3 across all tasks. Our most significant and consistent metrics across all tasks are the Shannon entropy, the duration, the mean of the speed and accelerations for point tasks, and the variations of angular speed and acceleration for the drawing and writing tasks. We show the box plots for some of those metrics in Figure 10.
Shannon entropy is a measure of the uncertainty in a signal, showing the erratic aspect of one’s handwriting, which is a marker of indecision for cognitively impaired participants. The differences measured in linear speed and linear acceleration, as well as duration, all show how significantly slower AD participants are from the control group. Finally, the differences in angular speed and angular acceleration are indicators of the swiftness of the writing. For detecting AD, the most significant metrics were all for writing and drawing tasks, which are much more demanding cognitively than the rest of tasks and mainly consisting of tasks designed for AD assessment, as opposed to the point and spiral tasks.
When comparing the tasks Copy Text and Copy Read Text, the difference in speed and horizontal acceleration is significantly higher for AD participants than CTL participants ($p$-value$/leq$0.001 for both), which shows how the multi-tasking has a way higher effect on cognitively participants.
Table 3: $p$-values and AUC for the best 10 metrics for AD vs CTL comparison. The table is repeated to fit the page. Significant $p$-values are bolded, as well as AUC values over 0.75
AD vs CTL Copy Image Copy Text Draw Clock Freewrite
p-val AUC p-val AUC p-val AUC p-val AUC
in-air avg. angular acc. 0.006 0.700 $<$0.001 0.791 $<$0.001 0.761 $<$0.001 0.760
in-air shannon entropy $<$0.001 0.873 $<$0.001 0.887 $<$0.001 0.866 $<$0.001 0.769
in-air std angular acc. $<$0.001 0.737 $<$0.001 0.818 $<$0.001 0.768 $<$0.001 0.821
in-air std angular speed $<$0.001 0.730 $<$0.001 0.826 $<$0.001 0.799 $<$0.001 0.774
on-tablet avg. hor. acc. $<$0.001 0.732 0.205 0.605 $<$0.001 0.788 0.003 0.698
AD vs CTL Numbers Point Dominant Point Non-Dominant Point Sustained
p-val AUC p-val AUC p-val AUC p-val AUC
in-air avg. angular acc. 0.025 0.691 $<$0.001 0.726 0.077 0.627 $<$0.001 0.740
in-air avg. acc. 0.685 0.536 $<$0.001 0.792 $<$0.001 0.850 $<$0.001 0.809
in-air avg. hor. acc. 0.357 0.593 0.021 0.663 0.006 0.688 0.007 0.698
in-air avg. speed 0.922 0.507 $<$0.001 0.797 $<$0.001 0.857 $<$0.001 0.827
in-air shannon entropy $<$0.001 0.794 $<$0.001 0.865 $<$0.001 0.810 $<$0.001 0.872
in-air std angular acc. 0.014 0.727 $<$0.001 0.804 0.003 0.708 $<$0.001 0.789
in-air std acc. 0.346 0.602 $<$0.001 0.716 $<$0.001 0.766 0.045 0.655
in-air std angular speed 0.025 0.694 0.021 0.665 0.131 0.608 0.112 0.615
in-air std speed 0.357 0.593 0.025 0.650 $<$0.001 0.737 0.052 0.647
Figure 10: Boxplots of the most significant metrics (AD vs CTL) presented across drawing and writing tasks for all groups of participants.
4.1.2 PD vs CTL
Over the 54 task-agnostic metrics, 19 are significant ($p\leq 0.05$) in the comparison between PD and CTL participants for at least one task, and 4 have p-values under 0.001. Table 4 shows the top 10 metrics with the lowest p-values across all tasks for the PD vs. CTL comparison.
Table 4: $p$-values and AUC for the best 10 metrics for PD vs CTL comparison. The table is repeated to fit the page. Significant p-values are bolded, as well as AUC values over 0.75.
PD vs CTL All Drawings Copy Image Copy Image Memory
p-val AUC p-val AUC p-val AUC
in-air avg. angular acc. 0.038 0.634 0.289 0.636 0.301 0.621
in-air avg. acc. 0.038 0.632 0.289 0.647 0.301 0.622
in-air avg. speed 0.038 0.622 0.289 0.635 0.359 0.610
in-air std angular acc. 0.009 0.670 0.227 0.679 0.301 0.659
in-air std angular speed 0.038 0.628 0.321 0.611 0.301 0.633
PD vs CTL Point Dominant Point Non-Dominant All Points
p-val AUC p-val AUC p-val AUC
in-air avg. angular acc. 0.002 0.723 $<$0.001 0.768 $<$0.001 0.723
in-air avg. acc. 0.189 0.611 0.116 0.624 0.007 0.611
in-air avg. hor. acc. 0.157 0.628 0.011 0.694 $<$0.001 0.637
in-air avg. vert. acc. 0.022 0.676 0.067 0.647 $<$0.001 0.650
in-air avg. angular speed $<$0.001 0.743 $<$0.001 0.768 $<$0.001 0.742
in-air avg. speed 0.169 0.620 0.106 0.630 0.008 0.608
in-air avg. hor. speed 0.169 0.616 0.019 0.678 $<$0.001 0.639
in-air avg. vert. speed 0.004 0.711 0.019 0.678 $<$0.001 0.682
in-air std angular acc. $<$0.001 0.738 $<$0.001 0.770 $<$0.001 0.723
in-air std angular speed $<$0.001 0.756 $<$0.001 0.759 $<$0.001 0.733
These metrics primarily revolve around in-air dynamics, variations, and averages of speeds and angular speeds. While notably discriminatory for point tasks, they also exhibit relevance, albeit to a lesser extent, for writing tasks like Copy Read Text and Freewrite. Additionally, we present box plots for these significant metrics in Figure 11 for further visualization and analysis.
Figure 11: Boxplots of some the best metrics extracted for PD vs CTL, all from angular speed and accelerations, presented for drawing and point tasks for all groups of participants.
For detecting PD, the most significant metrics were for drawing and point tasks. The point tasks were motor exercises designed for PD assessment, to measure PD’s typical tremor and bradykinesia. We can observe in these tasks that people with PD show significantly lower angular speed and angular acceleration in air, which could be indicators of hypometria [13]. The drawing tasks were much more complex tasks, mainly designed for cognitive evaluation but that still require basic motor skills to complete. In those tasks, the in-air angular speed is still a discriminative metric, but the main set of the metrics is constituted by the higher variations (as standard deviation or amplitudes) of the speed and acceleration on the tablet, a sign of tremor, typical of PD [12]. However, none of the task-agnostic metrics measured on the spirals tasks, mainly motor tasks often used for PD assessment, were discriminative enough.
4.1.3 PD vs PDM
Among the 54 task-agnostic metrics, 30 are significant ($p\leq 0.05$) in comparing PD and PDM participants for at least one task. We present the 10 metrics having the lowest p-values in Table 5 across all tasks. We only find significant metrics for Point and Spiral tasks here.
To distinguish PD from other parkinsonism, significant variations can be found in the Point and Spiral tasks and motor exercises designed for PD assessment. We observe few significant metrics for the Point tasks, mostly amplitudes of speed and acceleration, the best being for vertical velocity. PD participants have lower variations in their speed than PDM participants, which are still lower than healthy controls, as we can see in Figure 12. This means we observe less tremor in PDM participants than PD participants. The Spiral tasks reveal a difference in speed between PD and PDM participants, the latest being slower, as shown in Figure 13. This also makes the duration of the spiral drawing longer for PDM participants than for others.
Table 5: $p$-values and AUC for the best 10 metrics for PD vs PDM comparison. The table is repeated to fit the page. Significant p-values are bolded, as well as AUC values over 0.75.
PD vs PDM Spiral Dominant Spiral Pataka All Spirals
p-val AUC p-val AUC p-val AUC
on-tablet duration 0.323 0.680 0.128 0.686 0.003 0.685
on-tablet avg. angular acc. 0.323 0.696 0.128 0.726 0.003 0.687
on-tablet avg. hor. acc. 0.323 0.662 0.128 0.721 0.005 0.676
on-tablet avg. vert. acc. 0.323 0.662 0.128 0.710 0.003 0.685
on-tablet avg. hor. speed 0.323 0.676 0.128 0.736 0.003 0.696
on-tablet avg. vert. speed 0.323 0.700 0.128 0.752 $<$0.001 0.724
on-tablet shannon entropy 0.323 0.662 0.128 0.698 0.006 0.671
on-tablet std vert. speed 0.398 0.649 0.128 0.719 0.003 0.685
PD vs PDM Point Dominant Point Sustained All Points
p-val AUC p-val AUC p-val AUC
in-air amp vert. acc. 0.475 0.626 0.042 0.800 0.004 0.701
in-air amp vert. speed 0.540 0.607 0.063 0.767 0.005 0.690
Figure 12: Boxplots of some the best metrics extracted for PD vs PDM on the point tasks.
Figure 13: Boxplots of some the best metrics extracted for PD vs PDM for all the spiral tasks combined.
4.2 Task-Specific Results
After presenting the results of the task-agnostic metrics, we focus on the task-specific ones, which are crafted especially for certain tasks or groups of tasks.
4.2.1 Point Tasks Results
Figure 14 shows all metrics with their significance. None of the metrics are consistently significant across all tasks, but the radius average and radius standard deviation metrics exhibit the same behavior across all tasks: metrics from AD participants are higher than in the other groups. We observe a slightly higher radius average for PD participants compared to CTL, only significant when considering all point tasks, showing that the Point tasks capture some of the expected tremors.
Figure 14: Boxplots of the best point-specific metrics for all groups of participants.
4.2.2 Spiral Tasks Results
Figure 15 shows the boxplots of the spiral-specific metrics that were significant for at least one of the tasks. All the metrics exhibit a similar behavior across all tasks: PDM and AD participants take more time to draw their spirals than PD and CTL, AD and CTL participants draw more loops, which gives a lower time per loop (Fundamental period of the spiral) for PDM and AD participants. As participants were not instructed to produce a specific number of loops, it is hard to interpret if those findings are related to hyperkinesia in PD participants, finishing faster than the others, or if the decision to draw fewer loops and draw them slower could be linked to a form of cognitive decline.
Table 6: $p$-values and AUC of the specific metrics for all spiral tasks significant for at least one comparison. Significant p-values and AUC over 0.75 are bolded.
Spiral Metrics AD vs CTL PD vs CTL PD vs PDM
p-val AUC p-val AUC p-val AUC
N loops 0.092 0.573 $<$0.001 0.604 0.303 0.538
All avg radius diff 0.141 0.557 0.019 0.568 0.945 0.554
Spirals Duration $<$0.001 0.695 0.104 0.543 0.003 0.623
Number of Loops 0.370 0.563 $<$0.001 0.615 0.209 0.542
Fundamental Period $<$0.001 0.694 0.109 0.545 0.004 0.615
Dominant Duration $<$0.001 0.695 0.496 0.543 0.155 0.623
Fundamental Period $<$0.001 0.694 0.474 0.545 0.185 0.615
Non-Dominant N loops 0.714 0.573 0.018 0.604 0.166 0.538
Duration 0.123 0.695 0.043 0.543 0.028 0.623
Number of Loops 0.674 0.563 0.031 0.615 0.241 0.542
Fundamental Period 0.100 0.694 0.051 0.545 0.035 0.615
Pataka N loops 0.150 0.573 0.012 0.604 0.429 0.538
Duration $<$0.001 0.695 0.700 0.543 0.105 0.623
Number of Loops 0.406 0.563 0.007 0.615 0.620 0.542
Fundamental Period $<$0.001 0.694 0.658 0.545 0.103 0.615
Figure 15: Boxplots of the best spiral-specific metrics, for all groups of participants.
4.2.3 Numbers Task Results
Table 7: $p$-values and AUC of the significant metrics specific to the numbers task. Significant p-values and AUC over 0.75 are bolded.
AD vs CTL PD vs CTL PD vs PDM
Metrics p-val AUC p-val AUC p-val AUC
time in in-air 0.003 0.690 0.073 0.612 0.978 0.502
avg ans time 0.033 0.636 0.340 0.560 0.366 0.577
change in ans width 0.203 0.581 0.021 0.644 0.116 0.634
change in ans pressure 0.030 0.639 0.577 0.535 0.596 0.545
change in time btwn ans 0.007 0.673 0.088 0.607 0.963 0.504
Table 7 shows the Numbers-specific metrics’ results, studying mostly the variations in time and dimensions between successive answers in the Numbers Task. Of the few metrics that show a significant difference between groups of participants, the time taken to answer is the first metric differentiating between AD and CTL participants, AD participants being slower. In a similar fashion, AD participants are significantly slower over time as compared to CTL participants, an effect of the increased fatigue over the various answers. Finally, we observe that both AD and PD participants tend to input more and more pressure over time compared to CTL participants, once again a potential fatigue indicator.
4.2.4 Writing Tasks Results
Figure 16: Boxplots of the best writing specific metrics for all groups of participants.
In this section, we present the results for the writing tasks’s specific metrics. As shown in Figure 16, AD participants take a much longer Average Time Between Writing than all other participants, across all tasks, revealing a longer reflexion time between words. CTL participants generally have lower variations in the pressure they use than ND participants, across all tasks, a potential indicator of more stability in steadiness in the writing. PD participants use less and less pressure as they write, in contrast to the AD and CTL participants, and they also use a larger area to write, significant only in the Copy Text Task, which correlates with previous findings about PD participants writing larger than others [1].
Table 8: $p$-values and AUC of the specific metrics for all significant writing tasks. Significant p-values and AUC over 0.75 are bolded.
Writing Metrics AD vs CTL PD vs CTL PD vs PDM
p-val AUC p-val AUC p-val AUC
Copy change in pressure 0.810 0.508 0.043 0.624 0.115 0.573
Read press std 0.057 0.693 0.028 0.646 0.951 0.533
Text avg writing slope 0.028 0.524 0.060 0.599 0.710 0.504
avg time btwn write 0.009 0.710 0.264 0.583 0.665 0.702
press std 0.069 0.693 0.029 0.646 0.836 0.533
Copy area 0.112 0.612 0.044 0.542 0.897 0.562
Text avg writing slope 0.723 0.524 0.032 0.599 0.276 0.504
avg time btwn write 0.003 0.710 0.762 0.583 0.233 0.702
Freewrite press std 0.005 0.693 0.056 0.646 0.718 0.533
avg time btwn write 0.003 0.710 0.276 0.583 0.028 0.702
4.3 Classification Results
This subsection discusses the classification results obtained from the different handwriting tasks analyzed in this study. The results are discussed by examining each group comparison individually. Specifically, the following six group comparisons are analyzed: AD vs. CTL, PD vs. PDM, AD vs. CTL, AD vs. PDM, PD vs. CTL, and CTL vs. PDM. The models’ performance is evaluated in terms of accuracy (ACC), F-1 score (F1), and AUC. Table 9 presents results obtained using the Bagging classifier. The Appendix includes comprehensive experimental results from the RF and MLP classifiers.
Table 9: Classification results for pairwise group comparisons, using the Bagging model. The table reports experimental results in terms of accuracy (ACC), F1-score (F1), and AUC for each comparison. The best results for each metric for each group comparison are indicated in bold, while the second-best are italicized.
Task AD vs PD PD vs PDM AD vs CTL AD vs PDM PD vs CTL PDM vs CTL
ACC F1 AUC ACC F1 AUC ACC F1 AUC ACC F1 AUC ACC F1 AUC ACC F1 AUC
Copy Cube 71 68 81 50 48 57 67 66 69 67 64 69 64 63 71 57 51 55
Draw Clock 86 86 90 56 55 61 87 87 88 79 77 71 60 60 54 67 53 39
Copy Image 91 91 91 67 53 46 82 82 86 78 77 81 48 48 54 56 52 35
Copy Image Memory 73 73 83 65 51 51 64 63 62 67 67 63 69 69 69 54 48 47
Copy Read Text 77 75 81 59 44 35 67 67 72 76 74 78 61 60 62 71 69 79
Copy Text 77 78 88 63 58 74 81 82 89 73 72 83 63 63 64 67 65 66
Freewrite 69 69 76 66 52 45 71 71 75 73 70 65 48 48 47 77 77 83
Numbers 69 69 72 57 49 45 68 68 75 63 57 53 55 55 53 65 51 42
Point Dominant 91 91 93 66 52 47 84 84 85 82 82 78 66 65 68 57 55 53
Point Non-Dominant 84 84 94 67 56 52 85 84 85 73 71 72 68 68 66 71 71 72
Point Sustained 89 89 92 67 54 59 80 80 89 83 83 79 62 62 69 65 65 59
Spiral Dominant 67 64 69 67 53 50 60 60 63 75 69 57 54 52 54 60 57 60
Spiral Non-Dominant 65 63 59 64 63 66 57 55 58 79 78 64 41 41 33 60 57 50
Spiral PaTaKa 70 68 71 73 71 51 64 61 56 73 68 68 60 60 61 77 76 69
When comparing AD against PD, the Copy Image and the Point Dominant tasks exhibit the best performance in terms of accuracy (91) and F1-Score (91). The Copy Image task involves visual perception and fine motor skills, particularly impaired in PD. Moreover, the Point Non-Dominant task reports the highest AUC score (94), indicating its efficacy in discriminating between the two conditions [36]. This task might exacerbate the difficulties encountered by PD subjects when performing the task, as they are required to use their non-dominant hand.
In the comparison between the PD and PDM group, the Spiral PaTaKa task stands out with the highest accuracy (73) and F1 score (71). Spiral drawing tasks are known for their sensitivity in assessing motor skills due to their complexity and demand for coordinated movements. Given the significant impact of PD on motor function, tasks like spiral drawing can provide valuable insights into motor impairment severity. The task reporting the highest AUC is Copy Text.
For the comparison between AD and CTL, the Draw Clock task achieves the highest accuracy (87) and F1 score (87). This task assesses the ability to interpret and depict clock hands, a skill often compromised in individuals with early-stage dementia. The high discriminability of this task could be attributed to cognitive impairment present in the AD group, as indicated by their low MoCA score. Additionally, the Copy Text and Point Sustained tasks yield the highest AUC, indicating their utility in distinguishing between AD and CTL groups.
When contrasting AD with PDM, the highest accuracy (83) and the F1 score (83) are achieved in the Point Sustained task. This task requires that subjects maintain a pen vertically above a point while sustaining a vowel, demanding sustained attention and concentration. Given the attentional deficits commonly observed in AD, this task’s effectiveness in discriminating between AD and CTL and PDM groups is plausible.
In comparing PD and CTL groups, the Copy Image Memory task displays the best performance with an accuracy and F1 score of 69. This task assesses memory function, executive function, and visuospatial skills, which are often impaired in individuals with PD compared to CTLs. Additionally, the Copy Cube task yields the highest AUC, indicating its effectiveness in this comparison.
Finally, when comparing the CTL and PDM groups, the Freewrite task emerges as the most effective, achieving the highest accuracy and F1 score of 77 and an AUC of 83. This task involves language production and cognitive flexibility, domains commonly affected in individuals with PDM compared to healthy controls. The cognitive impairment exhibited by the PDM group is further confirmed by their low MoCA score (see Table X).
5 Conclusion
This study presents a composite array of handwriting interpretable metrics and examines their efficacy in distinguishing between various NDs, including Alzheimer’s disease and Parkinson’s disease. To accomplish this, we introduce a novel dataset recorded from 113 participants at the Johns Hopkins Hospital. This dataset comprises 46 healthy controls and 73 individuals diagnosed with various neurodegenerative conditions. The dataset includes recordings of 14 handwriting tasks conducted on a tablet, complemented by synchronized speech recordings and eye-tracking data (which are not considered in this study), resulting in 1840 handwriting files. Our investigation encompasses diversified metrics, both task-agnostic and task-specific. We aim to assess their ability to effectively differentiate between different groups of participants across various NDs.
First, we extracted metrics agnostic of the task at hand, to measure dynamic properties of the handwriting, such as the hesitations and swiftness through the speed of writing, the noise of the signal through entropies, or the hesitations by time measures. Our findings demonstrate that task-agnostic metrics contribute significantly to the discrimination of NDs, both quantitatively and qualitatively. Quantitatively, a substantial portion of the derived task-agnostic metrics (53 out of 54) exhibit significant differences between at least two groups of participants. Qualitatively because multiple metrics are very discriminant between groups (p-values $\leq 10^{-7}$), consistent across tasks of the same category, and consistent with previous findings. However, mainly due to the limited sample of participants, not all expected handwriting signs associated to each of the ND at hand have been successfully observed through those task-agnostic metrics. For example, none of the metrics really grasped a significant micrographia [12] from PD participants.
Because of the obvious limits of the generalist approach, we also explored a task-specific approach, where metrics were designed to capture specific signs for each group of tasks.
For AD participants, the extracted metrics grasped higher variations for the Point tasks than the CTL participants, surely due to fatigue, a lower speed and a higher duration when drawing spirals, more variations of the pressure applied and more time between each word written when writing paragraphs, as well as when answering arythmetic questions.
For PD participants, we observe a sligthly higher radius than CTL participants for the point tasks, draw less loops of a spiral for the same duration than CTL, while being significantly slower than PDM participants when drawing spirals. When writing, both PD and PDM groups have more variations in pressure on all tasks than CTL participants.
However, the comparisons between tasks brought almost no significant difference between controls and ND participants, which will require further efforts to highlight.
We can still observe that while the PDM group is younger than the others, the extracted metrics behave between the PD and CTL groups most of the time, showing the proposed metrics exhibit a robustness to the age variations.
The use of the proposed set of metrics for binary classification yielded interesting results for all the groups of tasks. Both Drawing tasks and Point tasks are very discriminative for AD participants (up to 91% Accuracy against the PD group, up to 87% Accuracy against the CTL group). Meanwhile, the Spiral tasks are more effective in discriminating PDM participants, in particular the Spiral PaTaKa task that shows 73% accuracy against PD, 73% against AD and 77% against CTL. Across all tasks, AD discrimination against CTL is almost always more accurate than PD against CTL, which confirm our hypothesis that handwriting in general contains valuable informations for AD assessement and characterization. In general, the variations of performances within a task group are non-negligeable, which is supprising given that the differences between similar tasks explored in the section 3.5.2 did not show any significance, except for Copy Text vs Copy Read Text, which mean those differences should be studied further, using more generalist systems to tackle groups of tasks at a time.
This study presents the handwriting component of a newly developed multimodal and synchronized dataset designed for the digital characterization and assessment of NDs. A set of metrics, both generalizable and task-specific, were introduced to capture dynamic characteristics of handwriting across various tasks. The results demonstrate that several of these metrics effectively discriminate between different participant groups, maintain consistency across tasks with similar characteristics, and align with established indicators of their respective neurodegenerative conditions.
The limitations observed in our approach suggest that generic solutions may not be fully suited to this specific application. Future research will focus on developing handcrafted machine learning models, tailored and trained to better capture the complexities of handwriting associated with neurodegenerative disease characterization. Additionally, given the multimodal and synchronized nature of the presented dataset, future efforts will explore the potential benefits of integrating all available modalities to enhance performance on the task.