### v3

# Explainable Metrics for the Assessment of Neurodegenerative Diseases through Handwriting Analysis

Abstract
Motor dysfunction is a common sign of neurodegenerative diseases (NDs) such as Parkinson's disease (PD) and Alzheimer’s disease (AD), but may be difficult to detect, especially in the early stages. In this work, we examine the behavior of a wide array of explainable metrics extracted from the handwriting signals of 113 subjects performing multiple tasks on a digital tablet, as part of the Neurological Signals dataset. The aim is to measure their effectiveness in characterizing NDs, including AD and PD. To this end, task-agnostic and task-specific metrics are extracted from 14 distinct tasks. Subsequently, through statistical analysis and a series of classification experiments, we investigate which metrics provide greater discriminative power between NDs and healthy controls and amongst different NDs. Preliminary results indicate that the tasks at hand can all be effectively leveraged to distinguish between the considered set of NDs, specifically by measuring the stability, the speed of writing, the time spent not writing, and the pressure variations between groups from our handcrafted explainable metrics, which shows $p$-values lower than 0.0001 for multiple tasks. Using various binary classification algorithms on the computed metrics, we obtain up to 87% accuracy for the discrimination between AD and healthy controls (CTL), and up to 69% for the discrimination between PD and CTL.
Index Terms:
Handwriting Analysis, Alzheimer's Disease, Parkinson's Disease, Digital Biomarkers, Neurological Signals
I Introduction
Neurodegenerative diseases (NDs), including Parkinson’s disease (PD) and Alzheimer’s disease (AD), progressively impair cognitive, motor, and behavioral functions by affecting both overlapping and distinct brain regions. To monitor these dysfunctions, many studies have explored non-invasive techniques using wearable or digital sensors. Among them, handwriting stands out as a promising modality due to its intricate integration of cognitive, motor, and perceptual processes [1, 2]. Handwriting production recruits a distributed neural network across cortical and subcortical structures, including regions involved in learning, executive function, praxis, vigilance, and motor coordination—spanning the cerebral cortex, basal ganglia, and cerebellum [3]. Consequently, abnormalities in handwriting patterns can serve as sensitive digital biomarkers of disease presence and progression [4], with micrographia being a well-established early symptom of PD [5].
Both paper-based and digital handwriting tasks have been employed to evaluate ND-related impairments by analyzing the spatial and temporal characteristics of handwriting signals [4, 5, 1]. Historically, most studies have focused on PD, largely due to the overt motor symptoms associated with the disease, including micrographia [6, 7, 8, 9], tremor [10], and bradykinesia [10]—collectively referred to as PD dysgraphia [1]. Handwriting analysis has also been explored for AD detection [11], as the disease’s prevalence continues to rise [12, 13, 14].
Numerous handwriting datasets have been introduced to study NDs, capturing either static images or dynamic time-series signals. However, most datasets are restricted to a single disease type and a single data modality, thereby limiting their experimental scope. For PD: HandPD [15]: Includes 92 participants (74 PD, 18 CTL), with image-based recordings of spiral and meander tracing. Naive Bayes classification achieved up to 78.9% accuracy. NewHandPD [15]: Expanded the previous dataset by adding pressure, acceleration, and tilt data from 14 PD and 21 CTL participants. Using CNNs, the study reached 83.77% accuracy on spirals and 87.14% on meanders. PaHaW [6]: Collected position and pressure data from spirals and repeated text in 37 PD and 38 CTL subjects. Using explainable metrics and SVMs, it achieved 81.3% accuracy. ParkinsonHW [16]: Recorded dynamic handwriting from 62 PD and 15 CTL subjects via a digital tablet. Cascaded non-neural classifiers achieved 84.67% accuracy on spirals and 90.91% on stability tasks [17].
For AD: Early work by [14] evaluated handwriting signatures from 29 AD and 30 CTL subjects, achieving a 3% error rate with a decision tree model. The dataset introduced by [13] included 23 AD, 12 MCI, and 17 CTL participants performing drawing and writing tasks on a tablet. Sentence copying reached up to 80.4% classification accuracy across three groups. The DARWIN dataset [12], now a standard benchmark, includes 100 AD and 90 CTL participants engaged in copying, memory, and graphical tasks. Fusing non-neural classifiers trained on individual tasks yielded a top accuracy of 94.29%.
Handwriting tasks for ND assessment can be categorized by complexity and task structure: Simple motor tasks involve drawing basic shapes such as spirals or straight lines, or stability exercises, such as maintaining a fixed pen over a tablet. These evaluate tremor, trajectory smoothness, velocity, and acceleration [18]. Written language tasks involve producing words or sentences. On tablets, these tasks enable the study of in-air movement and motor planning between words [18]. In-air time reflects cognitive preparation [5], and since in-air and on-surface dynamics are statistically independent [19], both provide complementary information. Increased in-air duration may reflect hesitation or planning deficits. Complex visuoconstructive tasks include copying figures or drawing clocks. The clock drawing test is an established tool for evaluating visuospatial and executive function across several NDs [20, 21]. Copying tasks allow examination of written response fidelity in varied contexts [22].
In this work, we introduce a new dataset composed of digital handwriting samples from 103 participants, including individuals diagnosed with AD, Mild Cognitive Impairment (MCI), PD, and Related Disorders (a heterogeneous group we refer to as PDM, comprising atypical Parkinsonism and misdiagnosed cases). Participants completed 14 tablet-based handwriting tasks specifically designed to probe cognitive and motor deficits across these conditions. We extract a large set of explainable, quantitative metrics, referred to as handwriting metrics, from these tasks. These include both: Task-agnostic metrics, which capture global signal properties across all tasks, regardless of their nature, and Task-specific metrics, which target metrics tailored to particular task types. We then assess which metrics best distinguish between groups through statistical analyses and binary classification experiments. In particular, we compare ND subgroups (PD, AD, PDM) to healthy controls, and contrast PD and PDM to identify discriminative metrics within motor phenotypes.
The main contributions of this paper are as follows:
  • Introduction of a novel, multi-condition handwriting dataset for ND assessment.
  • Inclusion of multiple disease groups, including an active control group (PDM), expanding beyond the narrower focus of prior datasets.
  • Extension of handwriting analysis from PD to AD and MCI, which have received limited attention.
  • Design and evaluation of a broad set of task-agnostic and task-specific metrics that align with expected behavioral patterns and reveal significant group differences.
  • Novel exploration of inter-task comparisons to analyze behavioral shifts when task parameters change (e.g., drawing hand).
The sections of the manuscripts are organized as follows: Section II presents the data collection pipeline and the main pre-processing stages. Section III introduces the task-agnostic metrics extracted from all tasks and the task-specific metrics designed for specific groups of tasks. Then, various methods used to gauge the significance of our metrics in the various tasks analyzed are described. Finally, we introduce the methods used for training classifiers for NDs assessment from the presented explainable metrics. Section IV shows and discusses our statistical and classification results. Section V contains the main findings, discussions, limits and conclusions of the study.
II Materials
In this section, we outline the data collection and preprocessing methods and the tasks evaluated in this research.
II-A Data Collection
The authors of this study collected a dataset from 113 participants from a larger digital biomarker cohort, NeuroLogical Signals, which our group has previously reported [23, 24, 25, 26, 27, 28, 29]. These participants were either categorized as healthy (CTL) or diagnosed and treated for a ND by a subspecialist neurologist or geriatric psychiatrist from Johns Hopkins Hospital. Participants diagnosed with NDs were categorized into aforementioned groups based on disease specific clinical diagnostic criteria diagnosis:
  • AD participants were enrolled from a University AD Center of Excellence, treated by an expert geriatric psychiatrist with an AD diagnosis based upon 2011 criteria [30]. However, most of these subjects were enrolled in serum or CSF biomarker studies confirming their diagnosis reflecting later revisions to diagnostic criteria [31].
  • MCI participants were characterized by established criteria, not yet meeting multidomain impairment or disability consistent with AD [32].
  • PD participants met international recognized diagnostic criteria at a "clinically established" [33] level of certainty and were followed by a movement disorders neurologist.
  • PDM participants presented with one or more core symptoms of PD, meeting a "Probable PD" diagnosis but which was later revised based on evolving characteristics or supportive testing. This is a mixed, active comparator group includes patients later diagnosed with Dementia with Lewy Bodies, Fahr’s Disease, Ataxia, Corticobasal Syndrome, Corticobasal Degeneration, Cervical Dystonia, Gerstmann Schenker Schyuler Syndrome, Wilson Disease, MSA-Parkinsonian, and Essential Tremor. In this study, this group is considered an active control group against the PD group.
All participants signed informed consent, and the data collection was approved by the Johns Hopkins Medicine Institutional Review Board. Participants with PD continued their usual pharmacological treatment and took dopaminergic medication before the recording session. Handwriting samples were collected using a Wacom One 13 tablet [^1] https://estore.wacom.com/en-us/wacom-one-13-touch-dth134w0a.html with the associated pen. This setup enabled us to record the pressure applied by the pen on the tablet surface and its position both on the tablet and in the air when in close proximity. Note the recording session also included collection of saccadometric and synchronized speech, although these two modalities were not utilized in this study. Participants completed 14 written tasks per session, detailed in Section II-B, guided by a research assistant. We collected recordings for each participant for at least one session, while 21 participants were recorded up to three times, when available and interested, with each session spaced six months apart. The handwriting data collected is part of a larger, multimodal dataset that is yet to be published. The distribution of participants, with their age, gender, and total number of tasks recorded, is presented in Table I Participants classified as AD and MCI were combined into a single group, named AD in this study. Only a subset of the recorded participants were kept in that study, to balance as much as possible the age and gender of the participants. However, given the low number of female participants, gender balancing for the AD group was not possible, so we chose to keep the whole group.
TABLE I: Distribution of the participants recorded in the dataset.
Category Number of % Females Age Files MoCA [34] UPDRS3 [35]
Participants recorded
CTL 42 57.1% 69y ($\pm 11$) 757 25.4
PD 35 40.0% 67y ($\pm 9$) 608 25.9 24.3
PDM 15 46.7% 53y ($\pm 14$) 283 25.8 21.5
AD 21 19.0% 70y ($\pm 7$) 745 19.7
Total 113 43.4% 66y ($\pm 12$) 1840
II-B Handwriting Tasks
In each session, we collected data from 14 handwriting tasks per participant. Due to myriad reasons, technical or human, the recordings of some tasks can fail, leading to a variable number of participants for each task. Participants having one or more missing tasks are still included in the study. Within this section, we provide a comprehensive description of tasks, their origin (what each of those tasks were used for, initially), and outline the experimental hypotheses we formulated for each. Tasks such as Point and Spiral tasks involved participants performing stability exercises three times: alternating hands, then with their dominant hand while concurrently speaking. Writing and Drawing tasks prompted participants to replicate an item from the screen, from memory, and generate new content. Participants were asked not to touch the tablet with their arm or hand while writing. Figure 1 illustrates a sample from each task type, encompassing participants from the CTL, AD, PD, and PDM groups.
Figure 1: Example of each type of task from all the four experimental groups. When on tablet, the color of the points transits from red to purple. When in air, the points are depicted in gray.
II-B1 Point Tasks
The point tasks are designed to assess the static motor control and stability of the participants. Each participant was asked to maintain the pen vertically above a point shown in the tablet without moving or touching the tablet for $10\text{\,}\sec$. More specifically, in the Point Dominant and Point Non-Dominant tasks, participants are asked to use their dominant and Non-Dominant hand respectively, and in the Point Sustained, they are asked to use their dominant hand, in addition, to sustain the phonation of the vowel A during the exercise.
II-B2 Spiral Tasks
The spiral tasks are also designed to assess the dynamic motor control and coordination of the participants. Several clinical rating scales [36, 37] require participants with PD to perform regular circular movement on paper or on tablet to assess their regularity or the presence of an eventual tremor [38, 39, 40]. The spiral test has also proven useful for AD assessment using a tablet [41]. The participants were asked to draw a spiral starting from the center of the tablet. More specifically, in the Spiral Dominant and Spiral Non-Dominant tasks, participants were asked to use their dominant and Non-Dominant hand respectively, and in the dual-task Spiral PaTaKa, they were asked to use their dominant hand, and simultaneously, to perform a spoken diadochokinetic task, i.e., repeat as fast as possible the syllables pa, ta, and ka in this exact order. However, our current focus lies solely on the handwriting data. We hypothesize that the simultaneous performance of tasks will make it more difficult for participants to use strategies to compensate for the motor impairments caused by NDs.
II-B3 Writing Tasks
The writing tasks were designed to assess the cognitive and motor capacities of the participants. Handwriting analysis has been used mostly for the assessment of PD, as some motor symptoms of the disease can be seen through some characteristic modifications of the handwriting, such as micrographia [42, 43]
Participants first engaged in the Copy Text task, wherein they were tasked with replicating a paragraph of text displayed on the screen. We then focus on the impact of multitasking by introducing the Copy Read Text task, where participants copied another text while simultaneously reading it aloud. Subsequently, participants' spontaneous writing abilities were captured through the Freewrite task, prompting them to write freely one or two sentences that must not contain any personal information. Lastly, we exploring handwriting a different context, the Numbers task, wherein participants were tasked with solving eight simple arithmetic operations summing two natural numbers not exceeding a total of 20, while saying the results of the sums out loud. It is worth noting that the Numbers task may be analyzed differently from the writing tasks due to its repetitive nature, necessitating a distinct analytical approach.
II-B4 Drawing Tasks
The Drawing tasks are also designed to assess the cognitive and visuo-constructive capacities of the participants.
First, in the Copy Image task, we asked participants to copy an image displayed on the screen. The image, is a complex figure in the style of those used in the Complex Figure Task [44]: asymmetric, non-patterned with internal and external details, as shown in the figure 2
Figure 2: Target image presented to the participants for the copy image and copy image memory tasks.
We further examined the impact of memory on this behavior by introducing the Copy Image Memory task. After copying the image in the Copy Image task, participants were then asked to reproduce it from memory approximately five minutes later, thus assessing visual memory and construction abilities.
Subsequently, participants were asked to perform two tasks: First, the Draw Clock task, which looked at their ability to draw all the required elements of a clock at an instructed time [45]. The Clock Draw Task is another established visuoconstructive screen studied in PD [21] and AD [34] which has been utilized to distinguish ND [46] and may be interpreted in a number of ways.
Then, the Draw Cube task, used in the MoCA test [34] and which has been proved to be also useful for PD assessment [47], examines their capacity to represent a three-dimensional object.
II-C Data Curation
Every task starts and concludes with a participant crossing an area marked as start or end on the screen, respectively, using a pen. However, the portion of the recording relevant to the analysis is shorter than the entirety captured. Hence, we manually eliminated the empty segments before and after the tasks commenced. Files lacking task-related responses are removed.
III Methods
This section outlines the metrics calculated for each task and the methods used to assess their significance among different experimental groups. We developed interpretable metrics from raw handwriting signals collected during various tasks. Our analysis has two aspects: first, we examine dynamic handwriting properties common to all tasks; second, we evaluate performance in specific tasks, such as stability in point tasks and proximity between drawing and target in copying tasks.
III-A Metrics computation pipeline
To compute the metrics analyzed in this study, we follow the pipeline described in Figure 3. Each data file can be described as a time series containing $N\in\mathbb{R}$ points:
  • The positions $X=(x_{i})_{i\in[[1,N]]}$ and $Y=(y_{i})_{i\in[[1,N]]}$
  • The time stamps $T=(t_{i})_{i\in[[1,N]]}$
  • The pressure $P=(p_{i})_{i\in[[1,N]]}$
We define a clean file as the list $\mathcal{F}=[X_{u},Y_{u},P_{u},T_{u}]$.
We computed various metrics from the clean files, targeting prioritizing motor, cognitive or behavioral elements. The metrics can be categorized relative to or irrespective of the task:
  1. 1.
    Task-agnostic (see Section III-B): these metrics measure primarily dynamic properties of the signal such as speed, noise, and basic ratios.
  2. 2.
    Task-specific (see Section III-C): these metrics are designed for specific tasks or groups of tasks.
III-B Task-Agnostic Metrics
Task-agnostic metrics quantify the dynamic aspects of the drawings. They are computed using several signal-processing techniques. The extraction process, illustrated in Figure 3, may be parsed:
  1. 1.
    computation of time series from the raw data, such as the angle between 2 consecutive points, the distance between two points, and multiple derivations (see Section III-B1).
  2. 2.
    segmentation of each series using only the segments gathered In-Air or On-Tablet, and additionally decomposing the On-Tablet sequence in a suite of strokes. A stroke being defined as the trajectory On-Tablet between two In-Air moments. The segmentation is described in Section III-B2
  3. 3.
    computation of statistics (e.g., mean, the standard deviation) on the time-series obtained from the derivation of time independent metrics, described in Section III-B3.
Figure 3: Pipeline used to extract task-agnostic metrics across all tasks. The amplitudes for angular speech and acceleration are not kept, which is why we have 54 and not 58 metrics.
III-B1 Time series
For every point of a given clean data file, we compute the angle and the $L_{2}$ norm between 2 consecutive points. Subsequently, we calculate the first and second derivatives for the following:
  1. 1.
    Angle: angular velocity and angular acceleration.
  2. 2.
    Norm: velocity and acceleration.
  3. 3.
    Positions: horizontal and vertical speed and acceleration.
  4. 4.
    Pressure: pressure variations.
These computations yield a total of 10 columns of data, which are utilized in subsequent analyses.
III-B2 Segmentation
We implemented segmentation to distinguish between segments occurring on the tablet surface and those in the air. Subsequently, we computed time-independent metrics for the entire file, focusing either solely on points recorded on the tablet or those recorded in the air. Given that the point task primarily involves actions in the air, statistical computations are exclusively performed on the in-air segments. Conversely, for spiral tasks, only segments registered on the tablet are utilized. However, for writing and drawing tasks, both in-air and on-tablet segments are utilized. Notably, metrics based on pressure are not retained for the in-air segments, as they would be derived from a constant value.
III-B3 Time-Independent Metrics
From the time series computed previously, we extract a range of statistics to characterize the dynamic properties of the signal, namely: the mean, the standard deviation and the amplitude of the signal.
We chose not to utilize the amplitude measure for angular velocity and accelerations, as it would always yield identical values. Furthermore, we calculated the Shannon Entropy based on the points' positions, whether on the tablet or in the air. To achieve this, we employ a Gaussian kernel density estimate (KDE) on the points' $X$ and $Y$ positions. This enabled us to determine the probability $q_{i}$ for each point $(x_{i},y_{i})$. Shannon Entropy is then defined as $H_{s}=-\sum_{i=1}^{N}q_{i}ln(q_{i})$. Following the computation of these metrics, we obtained a total of 54 task-agnostic metrics across the 14 tasks.
III-C Task-Specific Metrics
In this section, we detail the various task-specific metrics we computed. We divided the presentations of the metrics along the four main tasks' categories proposed in Section II-B, with the addition of a section specific to the Numbers task.
III-C1 Point Tasks Metrics
This section presents the metrics specific to the Point Tasks. Figure 4 depicts these metrics. Point Tasks are mainly used to measure the motor stability of participants and are mostly used for characterization of tremor, primarily relevant to PD or PDM (though tremor may also be observed in AD, due to copathology). Most dynamic metrics have already been explored as task-agnostic metrics, so we focused on the static metrics of the drawing: how symmetric and regular it is, and what the dispersion of the samples is. For this task, we calculated the geometric center of the pen's position over the digital tablet. We used this geometric center as a pivot to compare the distribution of the points with respect to the computed center, through 3 metrics:
  1. 1.
    Average Radius measures the dispersion of pen positions relative to the geometric center, providing a first measurement of stability that assesses the amplitude of variations for essential tremor.
  2. 2.
    Standard deviation of the Radius measures the variations in the distance from the center, a second measure more focus on dystonic tremors, for irregular vibrations.
  3. 3.
    Average Angle measures the average of the angle between a point and the center, which can be used as a proxy for the symmetry of the drawing. Tremor and bradykinesia in PD do not always affect all limb regions equally. An average angle >>>> 0 or <<0<<0 might implicate a strong tremor axis.
Figure 4: Schematic of the characteristics measured by different point-tasks specific metrics.
III-C2 Spiral Tasks Metrics
This section discusses static and dynamic metrics used to analyze hand-drawn spirals.
Those metrics are following three directions, illustrated in Figure 5:
  1. 1.
    Static characteristics of the ellipse bounding the spiral.
  2. 2.
    Static characteristics of the inside loops of the spiral.
  3. 3.
    Dynamic characteristics of the drawing, based on the time taken.
Figure 5: Spiral-specific metrics computation process
Ellipse characteristics
As shown in Figure 5, the participants produce rounded spirals for all the spiral tasks. Our approach involves initially characterizing the properties of the outer boundaries of the spiral as an ellipse, leveraging these properties as metrics. We used Principal Component Analysis (PCA) to find the directions of the semi-major and semi-minor axes of each spiral, then find the points further away from the geometric center of the spiral in both of those directions to obtain a measure of the semi-minor axis $a$ and the semi-major axis $b$. Using those values, we define a set of metrics:
  • Area of the ellipse: $Area=\pi*a*b$.
  • Eccentricity: $Ec=\sqrt{1+\frac{b^{2}}{a^{2}}}$. The eccentricity is a scalar $\in[0,1[$ that measures the compression of an ellipse when defined as a conic section. An eccentricity of 0 would be the one of a circle, while an eccentricity of 1 would indicate a parabola.
  • Ellipticity: $El=\frac{b}{a}$.
  • Degrees of rotation: The angle between the semi-major axis and the abscissa.
Spiral arms
Once the outside shape has been characterized, we looked into the properties of the arms of the spiral. Each spiral is split into segments, containing each one full loop. For each loop, we computed its radius as the average distances between its points and the geometric center of the spiral, which allowed us to compute another set of metrics:
  • Number of loops
  • The Average distance between loops: the difference between 2 loops radii is defined as the distance between those 2 loops. We take the mean of this measure for all pairs of consecutive loops.
  • The standard deviation of the loops radius differences: the standard deviation of the radius difference between consecutive loops. This metric quantifies the regularity of the spiral.
  • The slope of the loops radius differences: the slope of the radius difference between consecutive loops.
Dynamic analysis
Dynamic metrics are then computed based on the total time taken to complete the task:
  • Duration: total time taken to draw the spiral.
  • Time per loop: the average time taken to draw one loop.
  • Frequency: the average number of loops computed per second.
III-C3 Numbers Task Metrics
This section focuses on the metrics unique to the Numbers task. Participants engaged in solving addition problems by moving down the tablet, allowing for individual analysis of each response. This methodology enabled us to track the evolution of handwriting metrics as participants navigate through the task. For a comprehensive breakdown of these metrics, please refer to Figure 6 below.
Figure 6: Schematic of the metrics used for each answer of the Numbers task. The left graph shows the variations in pressures across the whole task, and the right schematic shows the metrics measured between 2 consecutive answers.
Average Width of Answer
The average width of each answer quantifies the size of the handwriting. It is computed by subtracting the minimum $X$ coordinate from the maximum $X$ coordinate for each answer and then averaging these values. People with PD have been found to have less control of their movement amplitudes and variance, resulting in compensatory longer strokes and larger handwriting under medication [48], as they are during the studied tasks.
Change in Width of Answer
The variation of width from one answer to the next. It is computed through a linear regression between the width of each answer and the question number. The metric itself represents the slope of this regression line. The variation in width is aimed to measure the presence of micrographia [42].
Change in Average Answer Pressure
The average answer pressure metric is a measure of the amount of pressure that a person applies while writing the answer. This metric can be leveraged as a time series metric to track its variation as the participant progresses through the task. It is computed through a linear regression between the average pressure of each answer and the question number. The metric itself represents the slope of this regression line. This metric might help pick up dystonia metrics, involuntary muscle contractions commonly seen in neurodegenerative diseases.
Time per Answer
The total duration metric measures the time a participant spends answering one of the questions. Given the absence of a predefined time limit for the Numbers task, this metric provides insight into the participant's pace in completing the task.
III-C4 Writing Tasks Metrics
In this section, we present the metrics extracted from the writing tasks: freewrite, copy text, and copy text memory[49]. The computation of these metrics is shown in Figure 7.
Figure 7: Schematic of the metrics used in writing tasks, illustrated for the words 'This is an example'.
Pressure Standard Deviation
The pressure standard deviation metric is a measure of the spread of the amount of pressure a person applies while writing. This measures how steady the hand is when writing by capturing the variations in pressure. Only the nonzero pressure values were considered when calculating this metric, as no pressure can be computed if the pen is not touching the tablet.
Average Time Between Writing
The average time between writing measures how long it takes for the participant to start writing after finishing a word. This shows how fluent the writing is and how stable the motor planning activity is. The metric was calculated by averaging the total time the pen was not touching the tablet.
Average Horizontal Distance Between Words
This metric was calculated by finding the difference between the $X$ value of when the pen is lifted and when it is put back down. Only the $X$ distance was considered because the $Y$ distance between the words is dependent on both the letters they are writing and how they write the letters.
Total Time
The total time metric gauges the duration required for task completion. It is determined by calculating the time elapsed from when the participant initially places the pen to start writing until they lift the pen after writing the last word. This metric holds greater significance for tasks like Copy Text and Copy Read Text, where participants' required text to write is the same.
III-D Statistical Analysis
After extracting a total of 76 metrics (i.e., 54 task-agnostic metrics and 22 task-specific metrics), we aim to evaluate their relevance for assessing and characterizing neurodegenerative diseases (NDs). Our objective is twofold: first, to identify which metrics effectively differentiate the CTL group from each of the disease groups (AD, PD, and PDM), and second, to assess whether handwriting metrics can distinguish PD from other forms of Parkinsonism (PDM).
To this end, we perform a Kruskal-Wallis H-test across the four experimental groups (CTL, AD, PD, PDM) for each metric, using the scipy.stats.kruskal function from the Scipy toolkit [^2] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.kruskal.html. For metrics showing a significant global effect (p<<0.05), we apply Dunn’s post hoc test to identify specific group differences. To control for false discoveries arising from multiple testing, we apply False Discovery Rate (FDR) correction using the statsmodels.stats.multitest.fdrcorrection function [^3] https://www.statsmodels.org/dev/generated/statsmodels.stats.multitest.fdrcorrection.html across the set of p-values grouped by task and test. Since we test 14 tasks across 4 pairwise group comparisons, we apply $4\times 14=56$ FDR corrections. To complete our analysis of the separability of those groups of participants, we measured the Area Under the Curve (AUC) for pairs of participant groups, which will be presented alongside FDR-corrected p-values.
Additionally, to investigate how individual metrics relate to cognitive and motor functioning, we compute the Pearson correlation coefficient between each of the 76 extracted features and both the MoCA [34] (Montreal Cognitive Assessment, for all participants) and UPDRS-III [35] (Unified Parkinson's Disease Rating Scale, motor section, only measured for the PD and PDM participants) scores across participants. To provide a more detailed analysis of motor function, we also compute correlations between each feature and the individual subdomains of the UPDRS-III score, including: rigidity, upper extremity bradykinesia, lower extremity bradykinesia, arising from a chair, gait, freezing of gait, postural stability, posture, gait and posture composite, kinetic tremor, postural tremor, resting tremor, global tremor score, and non-tremor components.
Significant results from both the group comparisons and correlation analyses are presented and discussed in the following sections.
III-D1 Grouped metrics
We hypothesized that certain metrics extracted for a specific type of task would exhibit consistent behavior across all tasks. To test this hypothesis, we grouped the tasks together. For each group, we conducted the same statistical comparisons outlined in section III-D. However, instead of analyzing the recordings from individual tasks, we considered all recordings within each task group. We then measured the significance of all metrics for each of these task groups as if they were a single task.
  • All spiral tasks: Spiral dominant, Spiral non-dominant, and Spiral PaTaKa.
  • All point tasks: Point dominant, Point non-dominant, and Point sustained.
  • All writing tasks: Freewrite, Copy Text, and Copy Read Text.
  • All drawing tasks: Draw Clock, Draw Cube, Copy Image, and Copy Image Memory.
Table II summarizes the tasks at hand sorted per group and the specific metrics proposed per group of tasks.
TABLE II: summary of the tasks at hand and their specific metrics.
Group Tasks Specific Metrics
Point Right Radius from centroid
Points Point Left Angle from Centroid
Point Sustained Std of the Radius
Area of the spiral
Spiral Right Eccenctricity
Ellipticity
Spirals Spiral Left Number of loops
Distance between loops
Spiral PaTaKa Time per loop
Frequency
Avg width of answer
Numbers Numbers Change in Avg pressure
Time per answer
Pressure Ratio
Copy Text Pressure std.
Writing Copy Text Memory Avg time between words
Freewrite Avg distance between words
Total time
Copy Image
Drawing Copy Image Memory CLIP similarity image-image
Draw Clock CLIP similarity text-image
Copy Cube
III-E Classification Experiments
In addition to conducting statistical analyses, we conduct binary classification experiments across various experimental groups using metrics extracted from each task. Namely, it involves a binary classification (e.g., AD or CTL) that facilitates the assignment of samples to their corresponding groups, thereby enabling the evaluation of the metrics' effectiveness in distinguishing between the categories under investigation. As for the statistical analysis, the feature set used in the classification models consists of both task-specific and general metrics, ensuring a comprehensive representation of handwriting dynamics. The number of metrics varies across tasks, with Copycube, Drawclock, Copymage Memory, and Copymage containing 54 metrics, while Copyreadtext, Copytext, and Freewrite include 62 metrics. The Numbers task has 68 metrics, whereas Point DOM and Point NONDOM each contain 28 metrics, and Point Sustained has 27 metrics. Additionally, Spiral DOM includes 38 metrics, while Spiral NONDOM and Spiral Pataka each have 37 metrics. These metrics capture various aspects of handwriting, including kinematics (e.g., velocity, acceleration, jerk), spatial properties (e.g., trajectory deviations, area), and pressure dynamics (e.g., mean pressure, pressure variability).
These experiments further evaluate the discriminative power of the metrics and the utility of tasks in extracting valuable information for characterizing different disorders. We are employing a range of classifiers, including Random Forest (RF), Bagging (BG), and Multi-Layer Perceptron (MLP)  [^4] To implement the different classifiers, we are using the Scikit-learn (Sklearn) library in Python. See https://www.tutorialspoint.com/scikit_learn/scikit_learn_introduction.htm. For each classifier, hyperparameter tuning was performed using nested cross-validation (NCV) to ensure optimal generalization performance. Specifically, for the BG classifier, we tuned max_samples (proportion of training samples used per estimator) and n_estimators (number of base estimators). For the MLP classifier, we optimized hidden_layer_sizes (number of neurons in the hidden layer). Finally, for the RF classifier, we tuned n_estimators (number of trees in the forest) and max_metrics (number of metrics considered per split). NCV is a widely used approach for assessing model performance, especially when dealing with limited data, as it helps provide an unbiased estimate of how well the model will generalize to new data. By incorporating two levels of data partitioning, NCV enables both model evaluation and hyperparameter tuning while minimizing overfitting and overly optimistic performance estimates [50, 51]. The NCV framework consists of two levels of cross-validation to ensure robust model evaluation and mitigate overfitting. In our implementation, the outer loop employs a 10-fold cross-validation, partitioning the dataset into 10 folds, with one fold held out for testing while the remaining nine are used for training. Within each outer training set, an inner 10-fold cross-validation is conducted to optimize hyperparameters. The best hyperparameter configuration is selected based on the average F1-score computed across all 10 inner folds, which helps mitigate variance in performance estimates and ensures stability in hyperparameter selection. The F1-score was chosen as the optimization criterion due to class imbalance. To prevent data leakage, cross-validation was strictly performed at the subject level. Speaker-independent folds were created to ensure that all samples from a given individual appeared only in either the training or test set within any iteration. This setup prevents data from the same subject from being in both training and testing, thereby avoiding overly optimistic performance estimates. The entire NCV process involved 100 hyperparameter tuning runs (10 outer folds $\times$ 10 inner folds), with final performance estimates derived by averaging results across the outer folds. This methodology provides a robust assessment of model generalization while maintaining strict speaker independence.
Figure 8: a) Machine learning pipeline was adopted for training and testing. b) Nested cross-validation for algorithm selection. In the graph, the nested cross-validation process is exemplified using 3 folds for training, one for testing in the outer loop, 2 for training, and one for validating in the inner loop. However, in our experiments, we used 9 folds for training, one for testing in the outer loop, 8 for training, and one for validating in the inner loop. This strategy was applied separately when training each of the classifiers adopted in the experiments.
IV Results
In this section, we first present the most significant metrics obtained by groups of metrics for each test realized. Then, we comment on the results obtained in the classification experiments for each task.
IV-A Task-agnostic Metrics Results
Out of 54 task-agnostic metrics computed, we found 53 were significant ($p\leq 0.05$) for, at least one task and one groups comparison. The complete table is available in the Appendix; however, we prioritize limited sets due to the considerable number of significant metrics. We present the metrics most effectively distinguishing AD and PD from control participants and separate PD from PDM participants. Subsequently, we will highlight metrics that exhibited robust behavior across multiple comparisons.
IV-A1 AD vs CTL
Over the 54 task-agnostic metrics, 50 are significant ($p\leq 0.05$) in comparing AD and CTL participants for at least one task, and 29 have p-values smaller than 0.001. We present the 10 metrics having the lowest p-values in Table III across all tasks. Our most significant and consistent metrics across all tasks are the Shannon entropy, the duration, the mean of the speed and accelerations for point tasks, and the variations of angular speed and acceleration for the drawing and writing tasks. We show the box plots for some of those metrics in Figure 9.
Shannon entropy is a measure of the uncertainty in a signal, showing the erratic aspect of one's handwriting, which is a marker of indecision for cognitively impaired participants. The differences measured in linear speed and linear acceleration, as well as duration, all show how significantly slower AD participants are from the control group, presumably due to processing inefficiency. Finally, the differences in angular speed and angular acceleration are indicators of the swiftness of the writing. For detecting AD, the most significant metrics were all for writing and drawing tasks, which are much more demanding cognitively than the rest of tasks and mainly consisting of tasks designed for AD assessment, as opposed to the point and spiral tasks.
When comparing the tasks Copy Text and Copy Read Text, the difference in speed and horizontal acceleration is significantly higher for AD participants than CTL participants ($p$-value $\leq$ 0.001 for both), which shows how the multi-tasking has a way higher effect on cognitive participants.
A limitation of our approach here is the lower number of participant for the AD/MCI category, which did not allowed us a separated analysis of the cognitive impairments against the participants presenting only AD's signs.
TABLE III: $p$-values and AUC for the best 10 metrics for AD vs CTL comparison. The table is repeated to fit the page. Significant $p$-values are bolded, as well as AUC values over 0.75
AD vs CTL Copy Image Copy Text Copy Read Text
p-val AUC p-val AUC p-val AUC
in-air avg. angular acc. 0.070 0.700 <<0.001 0.791 0.087 0.672
in-air shannon entropy <<0.001 0.873 <<0.001 0.887 <<0.001 0.827
in-air std angular acc. 0.019 0.737 <<0.001 0.818 0.034 0.702
in-air std angular speed 0.019 0.730 <<0.001 0.826 0.058 0.692
on-tablet avg. X acc. 0.017 0.732 1.000 0.605 <<0.001 0.761
on-tablet avg. Y acc. 0.078 0.680 1.000 0.597 <<0.001 0.754
AD vs CTL Numbers Draw Clock Freewrite
p-val AUC p-val AUC p-val AUC
in-air avg. angular acc. 0.125 0.691 <<0.001 0.761 <<0.001 0.760
in-air shannon entropy <<0.001 0.794 <<0.001 0.866 <<0.001 0.769
in-air std angular acc. 0.045 0.727 <<0.001 0.768 <<0.001 0.821
in-air std angular speed 0.125 0.694 <<0.001 0.799 <<0.001 0.774
on-tablet avg. X acc. 1.000 0.535 <<0.001 0.788 0.023 0.698
on-tablet avg. Y acc. 1.000 0.597 <<0.001 0.780 0.023 0.700
AD vs CTL Point Dominant Point Non Dominant Point Sustained
p-val AUC p-val AUC p-val AUC
in-air avg. angular acc. 0.016 0.726 1.000 0.627 0.030 0.740
in-air avg. acc. <<0.001 0.792 <<0.001 0.850 <<0.001 0.809
in-air avg. speed <<0.001 0.797 <<0.001 0.857 <<0.001 0.827
in-air shannon entropy <<0.001 0.865 <<0.001 0.810 <<0.001 0.872
in-air std angular acc. <<0.001 0.804 0.117 0.708 0.003 0.789
in-air std acc. 0.015 0.716 <<0.001 0.766 0.290 0.655
in-air std speed 0.159 0.650 0.004 0.737 0.390 0.647
Figure 9: Boxplots of the most significant metrics (AD vs CTL) presented across drawing and writing tasks for all groups of participants.
IV-A2 PD vs CTL
Over the 54 task-agnostic metrics, 10 are significant ($p\leq 0.05$) in the comparison between PD and CTL participants for at least one task, and 7 have p-values under 0.001. Table IV shows the top 10 metrics with the lowest p-values across all tasks for the PD vs. CTL comparison.
TABLE IV: $p$-values and AUC for the best 10 metrics for PD vs CTL comparison. The table is repeated to fit the page. Significant p-values are bolded, as well as AUC values over 0.75.
PD vs CTL All Points Point Dominant Point Non Dominant
p-val AUC p-val AUC p-val AUC
in-air avg. angular acc. <<0.001 0.723 0.041 0.723 0.009 0.768
in-air avg. acc. 0.027 0.611 1.000 0.611 0.443 0.624
in-air avg. X acc. <<0.001 0.637 0.659 0.628 0.032 0.694
in-air avg. Y acc. <<0.001 0.650 0.076 0.676 0.322 0.647
in-air avg. angular speed <<0.001 0.742 0.021 0.743 0.008 0.768
in-air avg. speed 0.031 0.608 1.000 0.620 0.437 0.630
in-air avg. X speed <<0.001 0.639 1.000 0.616 0.129 0.678
in-air avg. Y speed <<0.001 0.682 0.041 0.711 0.143 0.678
in-air std angular acc. <<0.001 0.723 0.052 0.738 0.010 0.770
in-air std angular speed <<0.001 0.733 0.021 0.756 0.009 0.759
These metrics primarily revolve around in-air dynamics, variations, and averages of speeds and angular speeds. They are notably discriminatory for point tasks. Additionally, we present box plots for these significant metrics in Figure 10 for further visualization and analysis.
Figure 10: Boxplots of some the best metrics extracted for PD vs CTL, all from angular speed and accelerations, presented for drawing and point tasks for all groups of participants.
For detecting PD, the most significant metrics were for drawing and point tasks. The point tasks were better suited to assess motoric metrics of PD, such as tremor and bradykinesia. Accordingly, we observed subjects with PD showed significantly lower angular speed and angular acceleration in air, which could be indicators of hypometria [52]. The drawing tasks were much more complex tasks, mainly designed for cognitive evaluation but that still require basic motor skills to complete. In those tasks, the in-air angular speed is still a discriminative metric, but metrics of variation (as standard deviation or amplitudes) of the speed and acceleration on the tablet were notable. These findings may represent, tremor, dystonia, dyskinesia, myoclonus other other complicating motor symptoms that develop in some persons with PD [10]. However, none of the task-agnostic metrics measured on the spirals tasks, mainly motor tasks often used for PD assessment, were discriminative enough.
IV-A3 PD vs PDM
Among the 54 task-agnostic metrics, only 6 are significant ($p\leq 0.05$) in comparing PD and PDM participants for at least one task. We only found significant metrics for the Spiral tasks here.
To distinguish PD from other parkinsonism, significant variations can only be found in the Spiral tasks, motor exercises designed for PD assessment. We observe few significant metrics for the Spiral tasks, mostly amplitudes of speed and acceleration, the best being for the vertical velocity. The Spiral tasks reveal a difference in speed between PD and PDM participants, the latest being slower, as shown in Figure 11.
Figure 11: Boxplots of some the best metrics extracted for PD vs PDM for all the spiral tasks combined.
Figure 12: Boxplots of the best point-specific metrics for all groups of participants.
IV-B Task-Specific Results
After presenting the results of the task-agnostic metrics, we focus on the task-specific ones, which are crafted especially for certain tasks or groups of tasks.
IV-B1 Point Tasks Results
Figure 12 shows all metrics with their significance. None of the metrics are consistently significant across all tasks, but the radius average and radius standard deviation metrics exhibit the same behavior across all tasks: metrics from AD participants are higher than in the other groups. We observe a slightly higher radius average for PD participants compared to CTL, only significant when considering all point tasks, showing that the Point tasks capture only little of the expected tremors.
IV-B2 Spiral Tasks Results
Figure 13 shows the boxplots of the spiral-specific metrics that were significant for at least one of the tasks. All the metrics exhibit a similar behavior across all tasks: PDM and AD participants take more time to draw their spirals than PD and CTL, AD and CTL participants draw more loops, which gives a lower time per loop (Fundamental period of the spiral) for PDM and AD participants. Subjects were not instructed to provide a minimum number of loops, therefore interpretation of task duration is difficult. The reduced number of loops in the PD cohort may have been an energetically rationale choice as the task is particularly difficult with persons with impaired dexterity. As PDM represents a mixed disease cohort, few conclusions relative to PD can be drawn. It is doubtful differences are due to cognitive impairment as overall MOCA performance was similar between these groups.
TABLE V: $p$-values and AUC of the specific metrics for all spiral tasks significant for at least one comparison. Significant p-values and AUC over 0.75 are bolded.
Spirals Metric AD vs CTL CTL vs PD PD vs PDM
p-val AUC p-val AUC p-val AUC
Spirals N loops 0.092 0.573 <<0.001 0.604 0.303 0.538
avg radius diff 0.141 0.557 0.019 0.568 0.945 0.554
Duration <<0.001 0.695 0.104 0.543 0.003 0.623
Number of Loops 0.370 0.563 <<0.001 0.615 0.209 0.542
Fundamental Period <<0.001 0.694 0.109 0.545 0.004 0.615
Dominant Duration <<0.001 0.695 0.496 0.543 0.155 0.623
Fundamental Period <<0.001 0.694 0.474 0.545 0.185 0.615
Non Dominant N loops 0.714 0.573 0.018 0.604 0.166 0.538
Duration 0.123 0.695 0.043 0.543 0.028 0.623
Number of Loops 0.674 0.563 0.031 0.615 0.241 0.542
Fundamental Period 0.100 0.694 0.051 0.545 0.035 0.615
Pataka N loops 0.150 0.573 0.012 0.604 0.429 0.538
Duration <<0.001 0.695 0.700 0.543 0.105 0.623
Number of Loops 0.406 0.563 0.007 0.615 0.620 0.542
Fundamental Period <<0.001 0.694 0.658 0.545 0.103 0.615
Figure 13: Boxplots of the best spiral-specific metrics, for all groups of participants.
IV-B3 Numbers Task Results
TABLE VI: $p$-values and AUC of the significant metrics specific to the numbers task. Significant p-values and AUC over 0.75 are bolded.
Metric AD vs CTL CTL vs PD PD vs PDM
p-val AUC p-val AUC p-val AUC
time in in-air 0.003 0.690 0.073 0.612 0.978 0.502
avg ans time 0.033 0.636 0.340 0.560 0.366 0.577
change in ans width 0.203 0.581 0.021 0.644 0.116 0.634
change in ans pressure 0.030 0.639 0.577 0.535 0.596 0.545
change in time btwn ans 0.007 0.673 0.088 0.607 0.963 0.504
Table VI shows the Numbers-specific metrics' results, studying mostly the variations in time and dimensions between successive answers in the Numbers Task. Of the few metrics that show a significant difference between groups of participants, the time taken to answer is the first metric differentiating between AD and CTL participants, AD participants being slower. In a similar fashion, AD participants are significantly slower over time as compared to CTL participants, an effect of the increased fatigue over the various answers. Finally, we observe that both AD and PD participants tend to input more and more pressure over time compared to CTL participants, once again a potential fatigue indicator.
IV-B4 Writing Tasks Results
Figure 14: Boxplots of the best writing specific metrics for all groups of participants.
In this section, we present the results for the writing tasks's specific metrics. As shown in Figure 14, AD participants take a much longer Average Time Between Writing than all other participants across all tasks, revealing a longer hesitation and intermitent pausing times, between words. CTL participants generally have lower variations in the pressure they use than ND participants, across all tasks, a potential indicator of more stability in steadiness in the writing. PD participants use less and less pressure as they write, in contrast to the AD and CTL participants, and they also use a larger area to write, significant only in the Copy Text Task, which correlates with previous findings about PD participants writing larger than others [1].
TABLE VII: $p$-values and AUC of the specific metrics for all significant writing tasks. Significant p-values and AUC over 0.75 are bolded.
Writing Metric AD vs CTL CTL vs PD PD vs PDM
p-val AUC p-val AUC p-val AUC
Copy Read Text press std 0.004 0.729 <<0.001 0.700 0.602 0.542
avg time btwn write 0.012 0.639 0.303 0.539 0.477 0.674
Copy Text press std 0.027 0.729 0.017 0.700 0.508 0.542
avg time btwn write 0.004 0.639 0.835 0.539 0.495 0.674
avg x dist btwn write 0.005 0.617 0.459 0.507 0.967 0.647
Freewrite press std <<0.001 0.729 <<0.001 0.700 0.628 0.542
avg time btwn write 0.019 0.639 0.539 0.539 0.046 0.674
IV-C Classification Results
This subsection discusses the classification results obtained from the different handwriting tasks analyzed in this study. The results are discussed by examining each group comparison individually. The models' performance is evaluated in terms of accuracy (ACC), F-1 score (F1), and AUC. Table VIII presents results obtained using the Bagging classifier. Since the BG classifier consistently outperformed the other two classifiers we tested—RF and MLP—in terms of accuracy, F1-score, and AUROC in the different tasks considered, we present and discuss the results obtained with this classifier in the main text for the sake of simplicity. The experimental results for RF and MLP are provided in the Appendix.
TABLE VIII: Classification results for pairwise group comparisons, using the Bagging model. The table reports experimental results in terms of accuracy (ACC), F1-score (F1), and AUC for each comparison. The best results for each metric for each group comparison are indicated in bold, while the second-best are italicized.
Task AD vs PD PD vs PDM AD vs CTL AD vs PDM PD vs CTL PDM vs CTL
ACC F1 AUC ACC F1 AUC ACC F1 AUC ACC F1 AUC ACC F1 AUC ACC F1 AUC
Copy Cube 71 68 81 50 48 57 67 66 69 67 64 69 64 63 71 57 51 55
Draw Clock 86 86 90 56 55 61 87 87 88 79 77 71 60 60 54 67 53 39
Copy Image 91 91 91 67 53 46 82 82 86 78 77 81 48 48 54 56 52 35
Copy Image Memory 73 73 83 65 51 51 64 63 62 67 67 63 69 69 69 54 48 47
Copy Read Text 77 75 81 59 44 35 67 67 72 76 74 78 61 60 62 71 69 79
Copy Text 77 78 88 63 58 74 81 82 89 73 72 83 63 63 64 67 65 66
Freewrite 69 69 76 66 52 45 71 71 75 73 70 65 48 48 47 77 77 83
Numbers 69 69 72 57 49 45 68 68 75 63 57 53 55 55 53 65 51 42
Point Dominant 91 91 93 66 52 47 84 84 85 82 82 78 66 65 68 57 55 53
Point Non-Dominant 84 84 94 67 56 52 85 84 85 73 71 72 68 68 66 71 71 72
Point Sustained 89 89 92 67 54 59 80 80 89 83 83 79 62 62 69 65 65 59
Spiral Dominant 67 64 69 67 53 50 60 60 63 75 69 57 54 52 54 60 57 60
Spiral Non-Dominant 65 63 59 64 63 66 57 55 58 79 78 64 41 41 33 60 57 50
Spiral PaTaKa 70 68 71 73 71 51 64 61 56 73 68 68 60 60 61 77 76 69
When comparing AD against PD, the Copy Image and the Point Dominant tasks exhibit the best performance in terms of accuracy (91) and F1-Score (91). The Copy Image task involves visual perception and fine motor skills, the latter of which is particularly impaired in PD. Moreover, the Point Non-Dominant task reports the highest AUC score (94), indicating its efficacy in discriminating between the two conditions [53]. This is intriguing as mirror movement or mirroring dystonia have been described in Task Specific Dystonia and other movement disorders as a way to unmask and characterize latent dystonia metrics [54]; in our PD cohort, this task may be abnormal in PD subjects who are prone to the development of dystonic metrics - a finding generally thought to be absent in AD.
In the comparison between the PD and PDM group, the Spiral PaTaKa task stands out with the highest accuracy (73) and F1 score (71). Spiral drawing tasks are known for their sensitivity in assessing motor skills due to their complexity and demand for coordinated movements. Given the significant impact of PD on motor function, tasks like spiral drawing can provide valuable insights into motor impairment severity. The task reporting the highest AUC is Copy Text.
For the comparison between AD and CTL, the Draw Clock task achieves the highest accuracy (87) and F1 score (87). This task assesses the ability to draw a clock with two hands correctly placed. This complex task includes a wide set of skills, including adequate function of spatial awareness, executive planning, error detection, motor coordination and execution as well as language processing and numerical processing. This task is often used to screen for intra-cortical and subcortical brain network deficiencies, observed notably in individuals with early-stage dementia. The high discriminability of this task could be attributed to cognitive impairment present in the AD group, as indicated by their lower MoCA score compared to other groups. Additionally, the Copy Text and Point Sustained tasks yield the highest AUC, indicating their utility in distinguishing between AD and CTL groups.
When contrasting AD with PDM, the highest accuracy (83) and the F1 score (83) are achieved in the Point Sustained task. Given the attentional deficits commonly observed in AD, this task's effectiveness in discriminating between AD and CTL and PDM groups is plausible.
In comparing PD and CTL groups, the Copy Image Memory task displays the best performance with an accuracy and F1 score of 69. This task assesses memory function, executive function, and visuospatial skills, which are often impaired in individuals with PD compared to CTLs. Additionally, the Copy Cube task yields the highest AUC, indicating its effectiveness in this comparison.
Finally, when comparing the CTL and PDM groups, the Freewrite task emerges as the most effective, achieving the highest accuracy and F1 score of 77 and an AUC of 83. This task involves language production and cognitive flexibility, domains commonly affected in individuals with PDM compared to healthy controls. The cognitive impairment exhibited by the PDM group is further confirmed by their low MoCA score (see Table X).
IV-D Correlation of the metrics with UPDRS-III and MoCa scores
To further explore the clinical relevance of the extracted handwriting metrics, we computed the Pearson correlation coefficient between each of the 54 task-agnostic metrics and the MoCA score [34], which evaluates global cognitive functioning, and the motor section of the Unified Parkinson's Disease Rating Scale (UPDRS-III) [35]. Only 13 out of 54 were significantly correlated with UPDRS-III scores ($p<0.05$), but none of them had $p$-values below 0.001, suggesting limited direct association between our handwriting metrics and motor symptom severity as assessed by UPDRS-III. However, a categorical examination of each subdomain of the UPDRS-III reveal a great variation of correlations depending on the domains, as show in Table IX. Asexpected, the tremor components are much more mesurable through handwriting than the postural and gait ones. The complete list of metrics with their individual significance is available in the appendix.
TABLE IX: Number of metrics over the 54 task-agnostic metrics produced being correlated with MoCa score, UPDRS-III score and its sub-components.
Score $p<0.05$ $p<0.01$ $p<0.001$ $p<0.0001$
MoCa [34] 33 22 15 9
UPDRS-III [35] 13 10 0 0
Rigidity 4 0 0 0
Upper bradykinesia 0 0 0 0
Lower bradykinesia 22 0 0 0
Arising from chair 2 0 0 0
Gait 20 4 0 0
Freeze of gait 8 2 2 2
Postural stability 0 0 0 0
Posture 1 1 0 0
Gait and posture composite 0 0 0 0
Kinetic tremor 19 5 0 0
Postural Tremor 15 6 0 0
Resting tremor 27 23 13 2
Global tremor 30 23 12 4
Non-tremor components 0 0 0 0
In contrast, a substantial number of metrics (33 out of 54) were significantly correlated with MoCA scores ($p<0.05$), indicating a strong relationship between handwriting behavior and cognitive performance. Among these, 9 metrics exhibited highly significant correlations with $p$-values below 0.0001, underscoring the potential of handwriting-based metrics as sensitive indicators of cognitive impairment.
V Discussion and Conclusion
The primary goal of this study was to explore the utility of handwriting as a digital biomarker for the assessment and characterization of neurodegenerative diseases (NDs), including Alzheimer's disease (AD), Parkinson’s disease (PD), and related Parkinsonian disorders (PDM). We introduced a novel dataset collected at Johns Hopkins Hospital from 113 participants (46 healthy controls and 73 ND patients) who completed 14 structured handwriting tasks on a digital tablet. This rich dataset comprises 1840 handwriting recordings and is part of a broader multimodal acquisition effort that also includes synchronized speech and eye-tracking data.
We proposed and evaluated a large array of interpretable handwriting metrics designed to capture motor, cognitive, and behavioral markers from handwriting signals. These metrics were categorized as either task-agnostic (general signal descriptors applicable across tasks) or task-specific (metrics tailored to the expected behavior in particular task types). Our aim was to assess the ability of these metrics to distinguish between ND subtypes and healthy controls using statistical analysis and binary classification experiments, while ensuring clinical interpretability and relevance.
V-A Task-Agnostic and Task-Specific Metrics: Findings and Limitations
Our analysis of task-agnostic metrics, those computed independently of the specific task performed, demonstrated that a substantial number (53 out of 54) significantly differentiated at least one group comparison (e.g., PD vs. control, AD vs. PDM). These metrics included speed, acceleration, angular velocity, entropy, and timing-related metrics. Notably, these general metrics were consistent across task categories and aligned with known ND symptomatology. For example, reduced speed and increased entropy were particularly salient in AD participants, aligning with executive dysfunction and impaired motor planning.
However, the task-agnostic approach also revealed limitations. While general metrics captured high-level differences, they occasionally failed to detect hallmark disease-specific signs. Most notably, micrographia, a classic symptom of PD, was not captured in a statistically significant way. This suggests that some ND-specific metrics may only manifest under conditions that tightly control for task type, hand dominance, or cognitive load.
To address this limitation, we introduced task-specific metrics, tailored to the intent and structure of each task. These revealed more nuanced and disease-relevant patterns. AD participants showed longer in-air times, slower execution, and greater pressure variability, particularly during cognitively demanding tasks such as paragraph writing and arithmetic. PD participants drew fewer spiral loops, moved more slowly during motor tasks, and exhibited greater pressure variability, overlapping with patterns observed in the PDM group. While the PDM group was younger, its behavior often fell between the PD and control groups, suggesting that our metrics are relatively robust to age differences.
Despite these strengths, cross-task comparisons (such as between dominant and non-dominant hand usage or between reading and non-reading conditions) showed limited discriminative power, with the exception of Copy Text versus Copy Read Text. These results suggest the need for more advanced methods to capture inter-task behavioral adaptations.
V-B Classification Performance and Observed Limitations
In this study, we prioritized classification using interpretable metrics, incorporating task-specific and general metrics across handwriting tasks. Our primary goal was to develop a model that provides meaningful insights into the motor and cognitive impairments associated with neurodegenerative diseases, ensuring that the extracted metrics remain clinically interpretable. This approach allows for a better understanding of the underlying mechanisms driving classification decisions, which is crucial for real-world applications in clinical settings. However, deep learning-based approaches, such as those leveraging raw handwriting images or spectrogram representations of handwriting signals, offer a promising alternative. In a recent study [55], we explored non-interpretable methods by applying CNN and CNN-BLSTM models to spectrogram representations of handwriting signals for neurodegenerative disease classification. Our results demonstrated that classification performance varied across handwriting tasks and spectrogram channel combinations, with CNN consistently outperforming CNN-BLSTM. While the current study focused on interpretable metrics, future work could investigate hybrid approaches integrating deep learning techniques with interpretable models to enhance classification performance while maintaining clinical transparency. This would allow for the benefits of deep learning’s feature extraction capabilities while ensuring that the results remain interpretable and useful for clinical decision-making.
V-C Clinical Applications and Implications
Handwriting analysis has strong potential as a non-invasive, low-cost, and accessible digital biomarker. It captures both motor and cognitive processes and can be easily deployed using commercial tablets or styluses. Clinically, such tools could support early screening, aid differential diagnosis, and monitor disease progression or response to treatment. The metrics developed in this study are interpretable and aligned with known clinical metrics, making them suitable for real-world use.
While handwriting analysis shows significant promise as a digital biomarker for neurodegenerative diseases, the current equipment setup remains a notable limitation for widespread clinical adoption. The data in this study were collected using a high-resolution graphics tablet capable of capturing position, pressure, and in-air movements with high precision. However, this hardware is relatively expensive and requires controlled conditions and trained personnel for operation. These factors restrict its applicability outside research environments and limit deployment in routine clinical settings or home-based assessments.
V-D Discussion
While the proposed metrics and classification results offer compelling evidence for the utility of handwriting in distinguishing neurodegenerative conditions, there are several limitations that affect generalizability. First, the models developed in this study are trained on a dataset collected under tightly controlled conditions with a homogeneous set of equipment and task protocols. Their applicability to new populations, clinical settings, or hardware setups remains to be validated. This lack of generalization could be addressed through the use of pre-trained systems, such as neural networks which can learn more transferable representations, or through the use of more diverse handwriting datasets.
Additionally, while handwriting captures a rich combination of motor and cognitive activity, it is only one window into neurodegeneration. The broader dataset collected in this study also includes speech and eye-tracking data, which have yet to be analyzed in combination with handwriting. Multimodal approaches are likely to improve both sensitivity and specificity by modeling interactions across domains such as language, motor planning, and attention.
Finally, extending the current dataset by increasing the number of participants, balancing demographic variables, and adding more task repetitions will improve statistical power and enable more reliable model development and evaluation.
V-E General Conclusion and Future Directions
This study introduced a diverse set of interpretable metrics for handwriting analysis and demonstrated their value in differentiating between neurodegenerative conditions using data from structured tablet-based tasks. The results validate handwriting as a meaningful digital biomarker and provide a strong foundation for future research and clinical translation.
Building on this work, our future efforts will focus on four main directions. First, we plan to increase the size and diversity of our dataset to improve model robustness and enable the study of more subtle effects, such as early-stage impairments and cross-condition overlap. Second, we will integrate the available speech and eye-tracking data to explore multimodal models that can capture interaction effects across cognitive and motor channels. Third, we will investigate neural network-based solutions, including the use of pre-trained models, to leverage broader datasets and improve generalizability across populations and devices. Finally, we will evaluate the performance of our models on more portable and affordable equipment, with the goal of developing lightweight and scalable tools suitable for deployment in real-world clinical and home settings.
Together, these future directions aim to bridge the gap between research prototypes and practical, interpretable, and accessible tools for early detection and monitoring of neurodegenerative diseases.
Appendix A Additional Tables
TABLE X: Number of files and participants for each task and each experimental group. The symbol $\Sigma$ denotes the total count across experimental groups for a given task.
Task # Files # Participants
(Code) CTL PD PDM AD $\Sigma$ CTL PD PDM AD $\Sigma$
Copy Cube (CC) 38 25 14 39 116 36 22 12 20 90
Copy Image (CI) 42 37 16 42 137 39 34 14 21 108
Copy Image Memory (CM) 41 33 15 21 110 38 30 13 13 94
Copy Read Text (CR) 42 28 16 42 128 39 25 13 21 98
Copy Text (CT) 42 34 15 44 135 39 31 13 20 103
Draw Clock (DC) 43 27 15 41 126 40 24 13 20 97
Freewrite (F) 43 35 16 46 140 41 32 14 21 108
Numbers (N) 45 35 17 34 131 41 33 15 19 108
Point Dominant (PD) 44 36 16 43 139 41 32 13 21 107
Point Non-Dominant (PN) 43 36 17 43 139 40 33 14 21 108
Point Left (PL) 44 36 17 43 140 41 33 14 21 109
Point Right (PR) 43 36 16 43 138 40 32 13 21 106
Point Sustained (PS) 40 34 15 40 129 38 31 13 21 103
Spiral Dominant (SD) 41 36 16 45 138 38 33 14 21 106
Spiral Non-Dominant (SN) 41 36 15 45 137 40 33 13 21 107
Spiral Left (SL) 43 36 16 45 140 40 33 14 21 108
Spiral Right (SR) 39 36 15 45 135 38 33 13 21 105
Spiral Pataka (SP) 43 32 16 44 135 40 29 14 20 103
Total 588 464 219 569 1840 42 35 15 21 113
TABLE XI: Correlation of all the task-agnostic metrics with the MoCa score, UPDRS-III score and all its sub-components. 'n.s.' Is used for non significant features, * for p-values<0.05<0.05, ** for p-values<0.01<0.01, *** for p-values<0.001<0.001, and **** for p-values<0.0001<0.0001.
MoCa UPDRS Rigidity Upper Lower Arising Gait Freeze Postural Posture Gait and Kinetic Postural Resting Global Non
 [34] III brady brady from of gait stability posture tremor tremor tremor tremor tremor
 [35] kinesia kinesia chair composite components
in-air duration **** ** n.s. n.s. n.s. n.s. ** * n.s. ** n.s. n.s. n.s. * ** n.s.
in-air avg. X speed *** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * ** *** **** n.s.
in-air std X speed n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * **** **** n.s.
in-air amp X speed * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. *** *** n.s.
in-air avg. X acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** *** *** n.s.
in-air std X acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** *** **** n.s.
in-air amp X acc. ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * * n.s.
in-air avg. Y speed *** n.s. n.s. n.s. n.s. * * n.s. n.s. n.s. n.s. * ** *** *** n.s.
in-air std Y speed n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. **** *** n.s.
in-air amp Y speed n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. *** *** n.s.
in-air avg. Y acc. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * ** *** *** n.s.
in-air std Y acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * *** **** n.s.
in-air amp Y acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. *** ** n.s.
in-air avg. speed ** n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** * *** *** n.s.
in-air std speed ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** n.s. ** * n.s.
in-air amp speed *** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** n.s. ** ** n.s.
in-air avg. acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** * *** *** n.s.
in-air std acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. ** * n.s.
in-air amp acc. ** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** n.s. ** ** n.s.
in-air std angular speed **** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. * n.s.
in-air avg. angular speed **** n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. * * n.s. * n.s.
in-air std angular acc. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s.
in-air avg. angular acc. *** n.s. * n.s. n.s. n.s. ** n.s. n.s. n.s. n.s. * * n.s. n.s. n.s.
in-air shannon entropy **** n.s. n.s. n.s. n.s. * * * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet duration **** ** n.s. n.s. n.s. n.s. ** * n.s. n.s. n.s. n.s. n.s. * ** n.s.
on-tablet avg. X speed ** n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std X speed * n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet amp X speed n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. X acc. n.s. n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet std X acc. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet amp X acc. *** n.s. n.s. n.s. n.s. n.s. n.s. **** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. Y speed * n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std Y speed * n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet amp Y speed n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. Y acc. n.s. n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std Y acc. n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet amp Y acc. ** n.s. n.s. n.s. * n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. speed n.s. n.s. * n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet std speed n.s. n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet amp speed n.s. * n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet avg. acc. n.s. n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet std acc. n.s. * n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. * n.s. n.s. n.s. n.s.
on-tablet amp acc. n.s. * n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std angular speed **** n.s. n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. * n.s.
on-tablet avg. angular speed **** ** * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
on-tablet std angular acc. **** ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. * ** ** n.s.
on-tablet avg. angular acc. *** ** n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. n.s. ** n.s. * n.s.
on-tablet avg. dP/dt ** n.s. n.s. n.s. * n.s. * n.s. n.s. n.s. n.s. n.s. n.s. * n.s. n.s.
on-tablet std dP/dt * ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. * n.s. ** ** n.s.
on-tablet amp dP/dt * ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** ** n.s.
on-tablet avg. d2P/dt2 * ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** ** n.s.
on-tablet std d2P/dt2 * ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. * n.s. ** ** n.s.
on-tablet amp d2P/dt2 * ** n.s. n.s. * n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s. ** ** n.s.
on-tablet shannon entropy **** n.s. n.s. n.s. n.s. n.s. ** **** n.s. n.s. n.s. n.s. n.s. n.s. n.s. n.s.
TABLE XII:
Task AD vs PD PD vs PDM AD vs CTL AD vs PDM PD vs CTL PDM vs CTL
ACC F1 AUC ACC F1 AUC ACC F1 AUC ACC F1 AUC ACC F1 AUC ACC F1 AUC
Random Forest
Copy Cube 64 63 73 53 52 58 63 64 70 67 65 69 64 63 66 65 62 55
Copy Image 83 83 89 53 46 62 79 79 84 78 78 80 52 52 49 58 55 64
Copy Image Memory 77 77 82 57 55 56 62 62 60 64 64 65 69 69 75 62 59 50
Copy Read Text 72 71 78 57 56 54 63 63 70 79 78 82 65 63 66 67 64 79
Copy Text 82 82 90 77 76 77 79 79 85 77 75 81 61 61 61 64 60 64
Draw Clock 78 77 88 50 49 52 77 78 88 79 78 68 60 60 56 64 64 67
Freewrite 73 73 73 59 56 70 71 71 77 79 79 78 59 58 55 75 74 79
Numbers 66 65 70 57 49 46 68 68 73 51 47 48 50 50 52 57 50 55
Point Dominant 88 88 93 43 41 39 83 82 84 82 82 80 62 62 62 59 57 53
Point Non-Dominant 88 88 95 65 64 56 84 83 89 78 76 73 68 68 71 67 66 67
Point Sustained 89 89 93 63 59 51 81 81 90 83 82 79 65 65 66 62 62 57
Spiral Dominant 72 71 71 60 55 56 68 68 68 73 70 58 49 48 56 69 67 64
Spiral Non-Dominant 63 62 63 68 67 68 70 69 63 75 74 67 45 43 41 64 62 59
Spiral PaTaKa 68 66 69 66 65 46 58 57 52 73 71 71 60 59 62 79 78 69
Multi Layer Perceptron
Copy Cube 78 78 81 53 51 52 56 56 57 69 69 69 57 57 59 52 53 51
Copy Image 84 84 90 67 63 56 72 72 82 69 68 67 59 59 62 49 48 48
Copy Image Memory 65 64 76 57 56 59 55 55 51 56 56 66 57 57 64 56 57 51
Copy Read Text 65 63 66 32 33 33 66 66 73 78 77 87 61 59 60 69 66 79
Copy Text 77 78 84 65 63 69 73 73 78 86 85 92 58 58 62 71 71 66
Draw Clock 73 72 79 59 59 51 79 79 88 77 76 69 54 54 61 52 53 50
Freewrite 60 59 63 59 56 50 62 62 72 75 75 77 45 45 44 73 72 75
Numbers 46 46 44 39 38 40 56 56 70 45 45 62 41 41 46 57 55 53
Point Dominant 88 88 92 57 53 57 81 81 84 74 71 75 62 62 73 64 62 65
Point Non-Dominant 83 83 89 57 55 57 84 83 82 75 72 72 59 58 68 64 64 65
Point Sustained 88 88 91 65 63 67 87 87 85 81 81 80 65 65 68 65 64 59
Spiral Dominant 67 66 74 47 42 38 59 58 58 78 75 51 56 56 55 74 74 64
Spiral Non-Dominant 61 59 60 75 73 59 69 68 69 73 66 70 48 48 49 57 51 52
Spiral PaTaKa 66 66 67 70 70 54 50 50 52 79 77 71 58 58 52 77 76 71